### v1

# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

Abstract
Monocular Re-Localization (MRL) is a critical component in numerous autonomous applications, which estimates 6 degree-of-freedom poses with regards to the scene map based on a single monocular image. In recent decades, significant progress has been made in the development of MRL techniques. Numerous landmark algorithms have accomplished extraordinary success in terms of localization accuracy and robustness against visual interference. In MRL research, scene maps are represented in various forms, and they determine how MRL methods work and even how MRL methods perform. However, to the best of our knowledge, existing surveys do not provide systematic reviews of MRL from the respective of map. This survey fills the gap by comprehensively reviewing MRL methods employing monocular cameras as main sensors, promoting further research. 1) We commence by delving into the problem definition of MRL and exploring current challenges, while also comparing ours with with previous published surveys. 2) MRL methods are then categorized into five classes according to the representation forms of utilized map, i.e., geo-tagged frames, visual landmarks, point clouds, and vectorized semantic map, and we review the milestone MRL works of each category. 3) To quantitatively and fairly compare MRL methods with various map, we also review some public datasets and provide the performances of some typical MRL methods. The strengths and weakness of different types of MRL methods are analyzed. 4) We finally introduce some topics of interest in this field and give personal opinions. This survey can serve as a valuable referenced materials for newcomers and researchers interested in MRL, and a continuously updated summary of this survey, including reviewed papers and datasets, is publicly available to the community at:
https://github.com/jinyummiao/map-in-mono-reloc
keywords:
Simultaneous Localization and Mapping, Positioning, Monocular Re-Localization, Pose Estimation
[…]
1 Introduction
Visual Re-Localization, as a typical state estimation problem, is a challenging and ongoing topic. Given a low-cost monocular camera, Monocular visual Re-Localization (MRL) tends to estimate 6 Degree-of-Freedom (DOF) poses, including orientation and position, with regards to scene map when the vehicle re-visits a mapped area VBL2018PR ; LCD2022TITS . This task holds great attentions in various autonomous applications, e.g., Virtual Reality (VR) Sven2014ECCV ; Jonathan2014TVCG , robots navigation Lim2015IJRR ; LATITUDE ; LocNeRF , Autonomous Driving (AD) TM3Loc ; MLVHM ; DA4AD ; RoadMap . Generally, MRL is commonly discussed to solve long-term association (so called loop closure) problem in Simultaneous Localization and Mapping (SLAM) system LCD2022TITS or recover pose as a state re-initialization in kidnapped robot scenarios Sotirios2020CVC .
Figure 1: An overview of the development in MRL. a) published papers titled by “visual re-localization” in each year, b) some of landmarks in MRL researches.
As the spatial model of the explored scene, map saves prior knowledge of scene (such as appearance, geometry, topology, etc.) and serves a reference coordinate for MRL. Generally, the MRL methods need to align current sensor data to scene map so that the ego pose can be solved, which can be consistently summarized as a two-step map-matching pipeline: build scene map in a specific form (mapping stage) and then localize based on the matching information between current data and the map (localization stage). Under the context of MRL, the data in the map is captured by sensors in the historical period, and it probably has extreme differences in visual conditions (like weather seqslam , illumination vprbench , day-night Asha2019ICRA , etc.) and even modality mono_i2p_vo ; cmrnet with the currently captured data, challenging the MRL solutions. Therefore, the representation form of data in scene map plays a vital role in MRL task and it affects the robustness of MRL methods against visual interference from real-world.
Figure 2: The structure of this survey.
Under such a context, various kinds of map have been involved in the existing MRL researches to achieve robust, efficient, and high-precision localization. In some proposals aiming at light-weight map, map is represented by frames with geographic location annotation (named as geo-tagged frame map) so that localization can be performed by image retrieval netvlad ; CoHOG or relative pose estimation arnold2022mapfree . For most of MRL methods tending to localization accuracy, map is built by visual point clouds with high-dimensional descriptors (named as visual landmark map) hloc1 ; hloc2 . In this formulation, the alignment can be performed by pixel-wise matching between local features orb ; superpoint . Some cross-modal MRL methods adopt raw point cloud map cmrnet ; mono_i2p_vo for its higher geometry accuracy and density. In the area of AD, the compact and vectorized semantic map, called vectorized High-Definition (HD) map, is widely utilized TM3Loc ; MLVHM . So the alignment process becomes matching between semantic instances. To achieve improvement by the benefit of End-to-End (E2E) way as other computer vision task alexnet ; faster-rcnn , some proposals learn to extract pose-related information from implicit map and localize itself, which takes current image as input and estimates poses posenet ; neumap . We believe that the scene map is implicitly represented in a form of network parameters and name such a kind of map as learnt implicit map. In this study, we find that MRL methods with the same kind of scene map usually perform in a theoretically similar way. Map determines the implementation way and even the localization performance of MRL methods.
However, existing surveys in related field do not discuss about the relationship between MRL methods and the scene map, which blocking the analysis and development of MRL. In the light of such a status, we aim to reduce the gap by reviewing existing MRL methods based on their utilized scene map. Specifically, as shown in Fig. 2, the MRL methods are categorized and introduced in five classes based on its used map, including not only traditional geo-tagged frame map (see Sec. 3), visual landmark map (see Sec. 4), and point cloud map (see Sec. 5), but also vectorized HD Map (see Sec. 6) and recently raised learnt implicit map (see Sec. 7). For each class, we fully review existing methods and deeply analyze the strength, weakness, evaluation metric, public benchmark. To boost the development of the community, we also discuss some opening problem in this field and provide some personal opinions.
Contribution The main contribution of this survey can be summarized as follows:
  • 1.
    To the best of our knowledge, this manuscript is the first survey exclusively paying attentions on visual localization with monocular camera as main sensor. We attempt to provide an in-depth review so that it can be used as a systematic reference material for newcomers and researchers acquainted with the MRL problem.
  • 2.
    The survey reviews MRL methods from a new perspective that we categorize the existing algorithms based on the representation form of utilized map. The relationships between map and MRL solution can be clearly studied.
  • 3.
    For each kind of MRL method with diverse scene map, we summarize the typical algorithms but also introduce the popularly used datasets, and evaluation metrics.
  • 4.
    Additionally, we also provide some opening discussions in MRL researches, which helps the community to further improve the MRL algorithms.
  • 5.
    As a final contribution, we build a continuously updated repository at github.com, including papers and datasets reviewed in this survey.
Organization The structure of this survey is as follows. Section 1 summarizes the contributions and paper structures. Section 2 describes the research background of MRL. From Sec. 3 to 6, the existing MRL methods are fully reviewed based on their scene map representation. The typical algorithms with various scene map will be evaluated in Sec. 8.2 and we analyze their advantages and disadvantages, respectively. Finally, Section 9 discusses some heated questions about MRL and Section 10 ends with conclusion and promising future directions.
Figure 3: Typical challenges of MRL methods. a) day-night changes with varying illumination tokyo247 , b) seasonal changes with different weather (sunny in summer and snowy in winter) seqslam , c) dramatic viewpoint changes university1652 , d) distinct places with visually similar objects, called perceptual aliasing problem malaga .
2 Background
2.1 Problem Formulation and Symbols Definition
We start by define MRL problem in theory and introduce the symbols used in this survey.
During the offline mapping stage, we build the scene map in a specific representation form. In those MRL method that scene map is explicitly represented and saved, scene map is $\mathcal{M}^{\mathcal{G}}=\left\{\textbf{m}^{\mathcal{G}}_{i}\right\}$, where $\textbf{m}^{\mathcal{G}}_{i}$ is the $i$-th represented item in the map, e.g., visual landmarks, point clouds, and semantics. Here, {}$\{\cdot\}^{\mathcal{G}}$ denotes the map is considered in a global coordinate of this scene $\mathcal{G}$ by default and can be simplified as =$\mathcal{M}=\{\textbf{m}_{i}\}$ if not specified. When a vehicle $v$ revisits such a mapped area, it captures the observation of the scene, that is, a monocular image $\mathcal{I}$ in MRL research. The vehicle $v$ can estimate its current ego-pose based on the current observations $\mathcal{I}$ given scene map information $\mathcal{M}$, which is defined as a MRL problem. Formally, the MRL problem can be written as
$\hat{\textbf{x}}^{\mathcal{G}}_{v}=\mbox{MRL}(\mathcal{I}~{}|~{}\mathcal{M})$ (1)
where ${}^^\hat{\{\cdot\}}$ denotes an estimated value, $\mbox{MRL}(\cdot)$ is a MRL solution, and $\textbf{x}^{\mathcal{G}}_{v}$ is the vehicle’s ego-pose with regard to $\mathcal{G}$.
Commonly, a 6 DOF pose $\textbf{x}^{\mathcal{G}}_{v}=\left[\textbf{R}^{\mathcal{G}}_{v}~{};~{}\textbf{t}^{\mathcal{G}}_{v}\right]$ is composed by a 3 DOF rotation $\textbf{R}^{\mathcal{G}}_{v}$ and a 3 DOF translation $\textbf{t}^{\mathcal{G}}_{v}$. The rotation can be represented as quaternion, rotation matrix, or Euler Angle. In some MRL application for ground AD vehicles, only 3 DOF pose is considered, i.e., lateral and longitudinal position $[x,y]$ and heading/yaw angle $\theta_{yaw}$.
For better readability, we define frequently used symbols here. The current image in localization stage (query image) is defined as $\mathcal{I}_{q}$ while a historical image at $t$ timestamp in mapping stage (reference image) is defined as $\mathcal{I}_{r(t)}$. Local features from $\mathcal{I}_{t}$ are denoted as $\mathcal{F}^{t}=\{\textbf{f}^{t}_{i}\}$ where a local feature $\textbf{f}^{t}_{i}$ contains a key point $\textbf{p}^{t}_{i}=[u,v]\in\mathbb{R}^{2}$ and corresponding local descriptor $\textbf{d}^{t}_{i}\in\mathbb{R}^{C}$ where $C$ is the dimension. A pair of matched local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r(t)}$ is denoted as $<\textbf{f}^{q}_{i},\textbf{f}^{r(t)}_{j}>$. The global feature of $\mathcal{I}_{t}$ is a global descriptor and denoted as $\mathcal{F}^{t}\in\mathbb{R}^{C}$. The items in $\mathcal{M}$ share the same symbols but in global coordinate, i.e., {}$\{\cdot\}^{\mathcal{G}}$.
2.2 Challenges of Monocular Re-Localization
MRL has been developed for decades and achieved many great success in many applications. However, tremendous challenges still block the development and usage of MRL methods in real word autonomy. Here, as shown in Fig. 3, some of primary challenges for MRL algorithms are given:
  • 1.
    Appearance change: Since monocular camera serves as the primary sensor in MRL solutions, the appearance changes of scene will dramatically affect the algorithms. The conditional appearance variance, e.g., weather seqslam , illumination vprbench ; tokyo247 , seasons seqslam ; 4seasons , and day-night changes benchmark6dof makes current image visually different from the map, thus challenging matching. And structural changes like dynamic occlusions slcd , and layout changes InLoc will interfere with the geometrical pose estimation.
  • 2.
    Viewpoint difference: When the camera viewpoint is quite different between current timestamp and mapping stage, MRL method will be struggled by limited co-visible area and changed layout of items. As typical instances, cross-view MRL for drones each_part_matters ; university1652 , opposite-view donot_look_back ; garg2018lost and even ground-to-aerial xview MRL for ground vehicles attract growing interests.
  • 3.
    Perceptual aliasing: In some scenaoris with visually similar or repeated textures, the MRL method will generate ambiguous estimation when distinct places have similar appearances, e.g., corridors rapnet ; openloris , and parking lots malaga , which is called perceptual aliasing problem.
  • 4.
    Generalization and scalability: The real world scenarios are infinite, and we cannot exhaust all the types, visual conditions, and interference of possibily occurred scenes. A practical real-world autonomy requires MRL methods to work stably in diverse scenarios, even in unseen environments, called generalization ability keetha2023anyloc ; SANet . Besides, the real world is unbounded, we also need MRL solutions scalable, which limits the unbearable increasement of map size and computational costs with vehicles continuously explore.
2.3 Comparison with other surveys
In decades of years, numerous reviews on autonomous robotics techniques have been raised, addressing various topics, e.g., place recognition LCD2022TITS ; Yin2022GeneralPR ; VPR_a_survey ; vpr_survey_ijcai ; VRP_a_survey_from_DL_perspective , ego and object localization VBL2018PR ; a_survey_on_visual_map_localization_using_LiDARS_and_camera ; visual_and_object_geo_localization_a_comprehensive_survey , SLAM past_present_future_SLAM ; a_survey_on_DL_for_SLAM ; visual_SLAM . As a milestone survey, Lowry et al. VPR_a_survey firstly introduced the concepts behind visual place recognition. Later, Garg et al. vpr_survey_ijcai refine the definition by introducing orientation limitation. With a growing number of visual place recognition methods based on deep learning, Zhang et al. VRP_a_survey_from_DL_perspective reviewed recently proposed solutions from the deep learning perspective. Yin et al. Yin2022GeneralPR presented general place recognition challenges. However, place recognition methods could only serve as a coarse localization that provide a rough pose approximation, which only cover few part of visual localization researches. In a_survey_on_visual_map_localization_using_LiDARS_and_camera , Elhousni et al. defined visual map localization as a two-stage process, namely, place recognition and metric map localization, and reviewed MRL methods using LiDAR, camera, or cross-modal sensors. Such a two-stage framework is only adopted by those MRL solutions using visual landmark map and point cloud map, and it cannot completely describe all the MRL solutions, such as newly emerged absolute pose regression-based methods posenet , scene coordinate regression-based methods neumap , etc.
By deeply analyze existing reviews about MRL methods, we find that although all these surveys have made great contribution to the community and boosted the progress of MRL researches, they have regrettably neglected the relationship between MRL methods and scene map, and thus cannot analyze MRL solutions in a unified framework. In this paper, we propose to review MRL methods from the perspective of scene map so that we can clearly categorize them based on the representation form of utilized map, thereby providing a comprehensive and deep review about MRL methods.
3 Geo-tagged Frame Map
Figure 4: A diagram of MRL methods using geo-tagged frames as map, including a) VPR and b) RPR methods.
During mapping stage, many frames are labeled by geographic position or even precise 6 DoF poses. These geo-tagged frames can be served as scene map in MRL methods. In this field, some proposals estimate ego-pose by historical pose of retrieved geo-tagged frames (Fig. 4 a) or estimated relative pose between query image and geo-tagged frame (Fig. 4 b).
3.1 Visual Place Recognition
Visual Place Recognition (VPR) algorithm aims to identify the re-observed places by retrieving reference frame when the vehicle goes back to a previously visited scene vpr_survey_ijcai ; VPR_a_survey , which is also adopted for loop closure detection in SLAM system LCD2022TITS . VPR regards the pose of retrieved reference frame as an initial or approximated pose of current query image. Formally, in VPR, the scene map $\mathcal{M}$ is represented by reference frames $\mathcal{I}_{r(t)}$ with pose $\textbf{x}^{\mathcal{G}}_{t}$, i.e., =$\mathcal{M}=\{\textbf{m}_{t}\}$ where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. Generally, given the query image $\mathcal{I}_{q}$, a VPR algorithm tends to retrieve the best-matched reference frame based on the image similarity:
$\hat{\textbf{m}}=\mathop{\arg\max}\limits_{\mathbf{m}}P(\mathcal{I}_{q},\mathcal{I}_{r(i)}~{}\mbox{of}~{}\textbf{m}_{i}|\textbf{m}_{i}\in\mathcal{M})$ (2)
where the matching function $P(A,B)$ provides the similarity or matching score between $A$ and $B$, and $\textbf{m}_{i}$ is a matching candidate, $\hat{\textbf{m}}$ is the best-matched one for $\mathcal{I}_{q}$. Then, the pose of current query image can be given by the pose of retrieved reference frame:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}\leftarrow\textbf{x}^{\mathcal{G}}_{\hat{\textbf{m}}}$ (3)
In early proposals, SeqSLAM seqslam and Fast-SeqSLAM fastslam direcly use down-sampled and normalized image patches to measure image similarity. However, raw image intensity is sensitive to visual interference. So VPR methods often extract high-level image features from images and measure feature similarity as a more reliable image similarity. In this feature-based scheme, image feature algorithm plays an important role in VPR solutions. In VPR, based on the receptive field, the image feature can be categorized into global feature hog ; vlad ; netvlad and local feature sift ; superpoint .
Global Feature-based VPR (GF-VPR): Global feature algorithm can directly represent the whole image as a compact matrix or vector so that the image similarity can be simply defined as a cosine distance $P(A,B)=({A\cdot B})/({\|{A}\|\|{B}\|})$ in Eqn. 2. Therefore, the core component of the GF-VPR methods should be the global feature algorithm. Before the deep learning stage, the global image features utilized in this area is designed to analyze the statistical distribution of image intensity, e.g., Gist Siagian , color histograms Ulrich , HOG hog . Zaffer et al. CoHOG improved HOG-based GF-VPR method by using image entropy to extract Regions-of-Interest and regional-convolutional descriptor matching. As a widely used solution, VLAD vlad converts $N$ local descriptors $\{\textbf{d}_{i}\}^{N}_{i=1}$ into a compact matrix $V$ using a codebook with $K$ clusters:
$V(j,k)=\sum^{N}_{i=1}a_{k}(x_{i})\left(\textbf{d}_{i}(j)-\textbf{c}_{k}(j)\right)$ (4)
where $\textbf{d}_{i}(j)$ and $\textbf{c}_{k}(j)$ are the $j$-th dimensions of the $i$-th descriptor and $k$-th cluster centre, respectively. $a_{k}(x_{i})$ denotes the membership of the $\textbf{d}_{i}$ to $\textbf{c}_{k}$, i.e., $a_{k}(\textbf{d}_{i})=1$ if $\textbf{c}_{k}$ is the closest cluster to $\textbf{d}_{i}$ and 0 otherwise.
Since deep learning-based methods dominate the visual tasks, some VPR methods have applied Deep Neural Network (DNN) to extract global features. As a milestone of DNN-based global feature, NetVLAD netvlad designs a trainable feature aggregation layer based on the principle of VLAD algorithm vlad . It replaces the non-differentiable VLAD pooling with soft-assignment of descriptors to multiple clusters, and construct the NetVLAD layer amenable to E2E training via back-propagation:
$V(j,k)=\sum^{N}_{i=1}\frac{e^{w^{T}_{k}\textbf{d}_{i}+b_{k}}}{\sum_{k^{\prime}}e^{w^{T}_{k^{\prime}}\textbf{d}_{i}+b_{k^{\prime}}}}\left(\textbf{d}_{i}(j)-\textbf{c}_{k}(j)\right)$ (5)
where $\{w_{k}\},\{b_{k}\}$, and $\{\textbf{c}_{k}\}$ are trainable for each cluster $\textbf{c}_{k}$.
Meanwhile, many other attempts are made by scholars to boost the representative ability of global feature. GeM GeM designs a trainable generalized-mean pooling layer that generalizes maximum and average pooling for feature aggregation. TransVPR transvpr could aggregate task-relevant features by self-attention mechanism in Transformers transformer . MixVPR mixvpr incorporates a global relationship between elements in each feature map in cascaded feature mixing modules. AnyLoc keetha2023anyloc leverages an off-the-shelf self-supervised foundation model features (such as DINOv2 oquab2023dinov2 ) with no VPR-specific training or finetuning. Although global feature can compactly represent images, the retrieval time of GF-VPR methods will linearly increase with the size of the database. And the robustness against viewpoint changes and dynamic occlusions of global features is also inferior than local feature DSFeat .
Figure 5: Four typical type of image representation in VPR researches, including a) image global feature, b) image global feature, c) image sequence feature, and d) image semantics.
Local Feature-based VPR (LF-VPR): A local feature algorithm extracts massive key points and corresponding descriptors from images. Then, the extracted local features are utilized to retrieve the best-matched reference frames. Traditional local features are hand-crafted based on human knowledge sift ; orb ; surf . Recently, many DNN-based local features have been adopted in VPR tasks. Early researches directly apply off-the-sheld DNN model pretrained by other computer vision tasks as local feature extraction module BoCNF ; PlaceRW ; convnet_pr . Then, specific DNN-based local features are introduced in LF-VPR researches and greatly improve the performance superpoint .
Directly matching between local features across two images is costly, local features need to be further processed to measure image similarity score. One common way is to aggregate extract local features to a compact global feature, e.g., VLAD vlad ; multivlad , ASMK ASMK , and BoW bow , and retrieve reference frames as GF-VPR methods. As typical solutions. DBoW dbow-iros ; dbow-tro extracts binary local features ORB orb from images and build vocabulary to aggregate local features so that an image can be compactly represented by the occurrences of visual words in the vocabulary (some general local features). Yue et al. yue1 ; yue2 improved it by introducing SuperPoint features superpoint . LRO LRO also utilizes SuperPoint features superpoint but it adopts an ASMK-based retrieval framework ASMK . In dynamic scene, local features located at dynamic objects/regions will confuse VPR methods, so Chen et al. slcd removed inconsistent dynamics by exploiting instance-level semantics. But semantics are hard to accurately judge the actual motion state of the objects, so DSFeat DSFeat learns to select stable and disciminative local features based on not only semantics but also attention information. These LF-VPR methods need to build offline vocabulary/codebook to aggregate local features, limiting the generalization ability. Some methods propose to build vocabularies in an online or incremental manner. In AVBoW , the authors proposed a novel updating strategy to automatically build and update vocabulary so that retrieval can still be performanced as dbow-iros ; dbow-tro .
Different from the solution aggregating local features and retrieving by aggregated feature similarity, some proposals tends to adopt probabilistic scheme in Eqn. 2 for finding matched reference frame. FAB-MAP fabmap and FAB-MAP 2.0 fabmap2 train a Chow-Liu tree to learn the co-visibility of features and then estimate the matching probability of candidates. IBuILD ibuild tracks BRISK features brisk across consecutive frames to incrementally generate vocabulary, and perform VPR by a likelihood function. iBoW-LCD ibow-lcd applies Bayesian framework by measuring the score of local features occurred in current image. LiPo-LCD lipolcd achieves further improvement by adding line-wise local feature in human-made environments. Tsintotas et al. AVWs dynamically segmented the incoming image stream to particular places and assigned places with visual words by using an on-line clustering algorithm . Then in BoTW , Tsintotas et al. improved the former work by tracking SURF features surf by KLT point tracker KLT to obtain robust tracked words and assigned them to the corresponding locations along the map, then VPR is performed by a binomial probability density function.
Figure 6: Three G-RPR solution proposed in arnold2022mapfree , namely, a) first solve scaleless relative pose by 2D-2D matches then recover scale by 3D-3D matches using estimated depth, b) first lift 2D-2D matches to 2D-3D matches by depth then estimate scaled pose by PnP+RANSAC solver, c) lift 2D-2D matches to 3D-3D matches using depth then estimate scaled pose.
Some methods exploit geometrical information from local features as a re-ranked matching score for matching candidates in Eqn. 2. Built upon NetVLAD netvlad , Patch-NetVLAD patch-netvlad re-orders candidates by cross matching of patch-level local DNN features with geometrical verification. As a de-facto standard in LF-VPR, researchers apply epipolar check across images and validate the matched reference by the number of feature inliers dbow-iros ; dbow-tro ; htmap ; AVWs ; BoTW ; fild2 . This method works properly in most scenes but fails if mismatched local features is dominated or pre-defined parametric model is incorrect. So, Yue et al. yue1 ; yue2 and Ma et al. LRO developed new methods to measure geometrical consistency. And Some other works choose to introduce temporal consistency in Eqn. 2 ibuild ; ibow-lcd ; AVWs but this scheme is limited to retrieval in sequential images. They follow a basic assumption that if $I_{q(t_{i})}$ and $I_{r(t_{j})}$ were matched, their adjecent images $I_{q(t_{i}-1)}$ and $I_{r(t_{j}-1)}$ should also be matched.
Sequence Feature-based VPR (SF-VPR): SeqNet seqnet and Delta Descriptor delta-des proposes to fuse temporal information that utilize the changes across image global features as sequence global feature. SeqVLAD SeqVLAD combines NetVLAD layer netvlad and sequence global features that exploits the temporal cues in a sequence. In these works, VPR first performs sequential matching across the image sequences then finds the most matched frame within the matched sequence.
Semantic-based VPR (S-VPR): Gawel et al. xview utilized semantic objects as features, and built a semantic topological graph based on the nearby relationship of semantic objects in images. The graph can be compactly represented by a global feature based on random walk deepwalk and S-VPR is performed in the GF-VPR way.
Summary: VPR algorithms aim to retrieve geographical nearby reference frame. Benefit from robustly representing images/map, current VPR solution can fully and precisely retrieve the frame. However, VPR perform localization by regarding the reference pose as an approximated query pose, which inherently limits the localization accuracy when the current trajectory is very different from the mapping trajectory. Thus, VPR usually serves as a coarse or initial localization step in practical localization system hloc1 ; hloc2 .
3.2 Relative Pose Estimation
Relative Pose Estimation (RPR) methods aims to estimate the relative pose between query image and reference image in the map. In this field, the scene map is also represented by images with pose, i.e., $M=\{\textbf{m}_{t}\}$, where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. The scene sometimes only contain one reference image $M=\textbf{m}_{r}=<\mathcal{I}_{r},\textbf{x}^{\mathcal{G}}_{r}>$ arnold2022mapfree . When the query image $\mathcal{I}_{q}$ is fed into the system, the RPR algorithm estimate the relative pose of $\mathcal{I}_{q}$ with regard to its (closest) reference image $\mathcal{I}_{r}$:
$\hat{\textbf{x}}^{r}_{q}=\mbox{RPR}(\mathcal{I}_{q},\mathcal{I}_{r})$ (6)
so that the absolute pose of $\mathcal{I}_{q}$ can be calculated by:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}\leftarrow\textbf{x}^{\mathcal{G}}_{r}*\hat{\textbf{x}}^{r}_{q}$ (7)
where *** is the pose transform operator. RPR methods avoid costly pre-scanning and reconstructing scene, enabling economical MRL in some new environment with extremely sparse records, which is commonly seen in Augmented Reality (AR) applications: User A shares a photo with its location in a new scene, any user B can instantly re-localize with regard to user A and perform interactions arnold2022mapfree . Generally, RPR methods can be categorized into geometry-based methods and learning-based methods. This section will review RPR algorithms from such a perspective.
Geometry-based RPR (G-RPR): The G-RPR methods perform feature matching between query image and reference image to obtain pixel-level correspondence, and then heuristically solve relative pose. According to multiple view geometry theory, the relative pose (relative rotation and scaleless translation) between two images with known camera intrinsics can be recovered by decomposing the essential matrix MVG . The essential matrix is classically estimated by matching local features, followed by robust estimation, such as using a 5-point solver 5points inside a RANSAC circulation ransac . Later proposals further improve such a basic formula by learning better features sift ; surf ; orb ; superpoint ; disk ; alike ; aliked ; r2d2 , better matching superglue ; LoFTR ; adalam ; lindenberger2023lightglue and better robust estimators ransac_analysis ; good_correspondence ; MAGSAC ; MAGSAC++ ; NGRANSAC ; ACNe .
In the early G-RPR solutions, scholars utilized scaleless pairwise relative poses between the query image and multiple reference images to triangulate the scaled, metric pose of query image EssNet ; ExReNet . However, in the scene where only one reference image exists, such a scheme fails to recover scale MVG . In arnold2022mapfree , as show in Fig. 6, authors proposed to utilized estimated monocular depth DPT ; planeRCNN to a) provide scale or b)lift 2D-2D correspondences to 2D-3D or c) 3D-3D correspondences, and concluded the second solution back-projecting reference image to 3D space achieves best performance, which is technically converting G-RPR to typical Perspective-n-Point (PnP) problem. And arnold2022mapfree also concludes better feature matching could boost the G-RPR methods.
Learning-based RPR (L-RPR): With the help of DNN model, L-RPR algorithms predict the relative pose from two input images in an E2E way and bypass explicit estimation of 2D correspondences. Some methods first estimate pose up to scale and then recover scale by triangulation CNNSPP ; ExReNet ; Zakaria2017ICCVW , while others aim to directly estimate metric relative pose DistillPose ; RelocNet ; RPNet . Both RelocNet RelocNet and ExReNet ExReNet show great generalization ability when perform RPR on unseen scenarios. Since L-RPR methods do not need to match between query image and reference image, they generally have low requirements of texture in the scene. Thus, L-RPR has been applied in scenarios that are challenging for G-RPR approaches. Cai et al. Cai2021ExtremeRot focused on estimating the relative rotation between two images in extreme cases, including when there is no overlap between the two images. Similarly, the method in DirectionNet estimates scaleless relative pose for pairs of images with very low overlap.
In arnold2022mapfree , authors combined two type of PRP methods and designed a network to predict 3D-3D correspondence between two images, adopting KeypointNet keypointnet . The relative pose is then solved as G-RPR method using Orthogonal Procrustes Procrustes inside a RANSAC loop ransac . The experimental results claimed that this solution outperforms the L-RPR method, indicating that involving multiple view geometry knowledge in neural network may benefit RPR algorithms.
Summary: RPR perform localize by estimating relative pose between query and reference images. Since scale recovery for pairwise RPR is not perfectly solved yet, RPR methods cannot achieve very high-precision localization performance. But, the low requirement of scene map still let RPR attract growing attention in light applications like AR.
4 Visual Landmark Map
As the most popularly applied representation format of scene map, visual landmarks is constructed by Structure from Motion (SfM) colmap or SLAM orbslam3 . Visual landmarks is some informative and representative 3D points that lifted from 2D pixels by 3D reconstruction, and they are associated with corresponding local features in various observed images including 2D key point and high-dimensional descriptor as shown in Fig. 7, i.e., scene map can formatted as =$\mathcal{M}=\{\textbf{m}_{i}\}$, where $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},\{\mathcal{F}^{t}_{i}\}_{t}>$ and $\mathcal{F}^{t}_{i}$ is a set of local features from reference image $\mathcal{I}_{t}$ associated with $\textbf{m}_{i}$, $\textbf{p}^{\mathcal{G}}_{i}$ denotes the 3D location of $\textbf{m}_{i}$ in global frame $\mathcal{G}$. Among the Visual Landmarks-based MRL methods (VL-MRL), Hierarchical Localization (HLoc) hloc1 ; hloc2 is the most famous framework and has been widely used in many applications. Up to now, it still dominates the long term visual localization task [^1] https://www.visuallocalization.net/benchmark/.
Figure 7: A diagram of MRL method with visual landmark map.
In HLoc framework hloc1 ; hloc2 , scene map $\mathcal{M}$ composed by 3D visual landmarks with descriptors is offline built via SfM reconstruction colmap . During online localization stage, query image $\mathcal{I}_{q}$ is matched with retrieved reference image $\mathcal{I}_{r}$ and the resulting 2D-2D matches are lifted to 2D-3D matches between $\mathcal{I}_{q}$ and scene map based on $\mathcal{M}$, which can be used to solve scaled pose as a typical PnP problem. For clarify, we split VL-MRL algorithms into two step based on their functions (Step 1.A and 1.B are alternative):
  • 1.A
    extract local feature from $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ (we assume only one $\mathcal{I}_{r}$ is utilized here):
    $\begin{split}\mathcal{F}^{q}=\{\textbf{f}^{q}_{i}\}_{i}&=F_{\mbox{feat}}(\mathcal{I}_{q})~{}~{}\mbox{during localization}\\ \mathcal{F}^{r}=\{\textbf{f}^{r}_{i}\}_{i}&=F_{\mbox{feat}}(\mathcal{I}_{r})~{}~{}\mbox{during mapping}\end{split}$ (8)
    Then, match local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ and obtain 2D-2D correspondences:
    $\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}=\mbox{match}(\mathcal{F}^{q},\mathcal{F}^{r})$ (9)
    where $i$ is the indices of matched local features.
  • 1.B
    or jointly extract and match local feature between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$:
    $\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}=F_{\mbox{fm}}(\mathcal{I}_{q},\mathcal{I}_{r})$ (10)
    where $i$ is the indices of matched local features.
  • 2
    get the associated visual landmarks of $\mathcal{F}^{r}$ in $\mathcal{M}$ so that 2D-2D correspondences can be lifted to 2D-3D correspondences, and the pose can be solved as a typical PnP problem:
    $\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{solvePnP}(\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}~{}|~{}\mathcal{M})\\ &=\mbox{solvePnP}(\{\textbf{f}^{q}_{i},\textbf{f}^{\mathcal{G}}_{i}\}_{i})\end{split}$ (11)
    where $i$ is the index of matched local features or visual landmarks.
In this section, we review related proposals in the two steps of VL-MRL methods, namely, local feature extraction then matching (Step 1.A), joint local feature extraction and matching (Step 1.B), and pose solver (Step 2). And we also review some proposals aiming at other perspectives in VL-MRL methods.
4.1 Local Feature Extraction-then-Matching
Local Feature Extraction: As previously mentioned in Sec. 3.1, local feature detects massive salient pixels in an image (denoted as key points), and describes the neighboring area of the key point using a high-dimensional vector (denoted as descriptor). Under the context of VL-MRL methods, local feature is utilized to obtain accurate 2D-2D pixel-wise correspondences between images. The key points detected by local feature algorithms should be repeatable under different image conditions and qualities while located in salient regions zhang2017learning so that visual landmarks built by local features can be sparse but informative to represent the whole scene rapnet ; DSFeat . And descriptors should be designated to enable accurate correspondence establishment between detected feature points across different views l2net ; hardnet ; sosnet ; geodesc ; contextdesc , so they need be invariant against typical visual interference. Additionally, local feature should be accurate.
The early local features are built upon the human prior knowledge about the quality of pixels. As milestones in this area, the float SIFT sift , SURF surf and binary BRISK brisk , BRIEF brief , ORB orb have been widely used and achieved great successes in many autonomous applications. However, due to the limited presentation ability, these hand-crafted local features are challenged by complex visual interference in real-world scenarios. Since entering the perception age ruled by DNN, researchers have begun to introduce deep learning technique in local feature extraction. As a beginning, SuperPoint superpoint utilizes a shared encoder and two heads to automatically detect key points and extract descriptors, respectively. It is first trained on synthetic samples and then fine-tuned in two images generated by the homographic adaptation. After that, the training datasets with posed images and 3D map are constantly proposed, the ground-truth correspondence between different views of the scene can be available, facilitating the training of DNN-based local features. D2-Net d2net is optimize by apply metric learning between two co-visible patches and achieve better performance, but it extracts local feature on a feature map with reduced resolution, so the accuracy of matching and localization is limited. ASLFeat aslfeat improves by detecting on feature map with original resolution. ALIKE alike and ALIKED aliked support sub-pixel-level detection. Thus, their accuracy are boosted with a lot of margin. As a commonly applied strategy, multiple scale (MS) detection can help local feature extraction. Some works explicitly apply MS detection d2net ; sekd , some other works implicitly apply MS detection in network by fusing multiple features from feature pyramid inherited by CNN model aslfeat ; alike . To make DNN-based local feature get rid of costly training data collection, SEKD sekd proposes a self-evolving framework to train CNN model without any annotations or pre-processing on the training data. Some researchers apply attention mechanism in CNN model to further enhance the representative ability of local features delf ; mda . These methods train CNN model as GF-VPR tasks (see Sec. 3.1) and then detect local features based on estimated attention mask of trained CNN model. To facilitate the interpretability of trained attention mask, DSFeat DSFeat further combine attention information and semantic information to select stable and discriminative local features. Different from the methods trained with metric learning scheme, DISK disk train CNN model by reinforcement learning scheme where the feature extraction and matching procedures can be trained in an E2E manner, thereby greatly boosting the performance of local features.
Figure 8: The processed Frame Per Second (FPS) and Mean Matching Accuracy (MMA) on HPatches dataset hpatches1 ; hpatches2 . The FPS metrics of SIFT sift , SURF surf , D2-Net d2net , SEKD sekd , GLFeat glfeat are evaluated on a NVIDIA RTX GPU glfeat , whereas LF-Net lfnet , SuperPoint superpoint , R2D2 r2d2 , ASLFeat aslfeat , DISK disk , ALIKE alike , and ALIKED aliked are evaluated on a NVIDIA GeForce RTX 2060 GPU aliked . “SS” denoted single scale detection while “MS” is multiple scale detection.
Limited by the locality of CNN, most existing local features description methods only learn local descriptors with local information and lack awareness of global and surrounding spatial context. Thus, MTLDesc MTLDesc and GLFeat glfeat aggregate non-local awareness into local feature extraction. Besides, as shown in Fig. 8, most of the state-of-the-art local features cannot meet the requirements of efficiency even with high-performance GPU devices, which blocks the usage of local feature algorithms in real-time applications. Zhao et al. specially designed light CNN model for real-time local features alike ; aliked . Thin neural networks usually sacrifice representative ability, so strong supervision for local feature training is required to achieve balance between matching accuracy and computational efficiency alike ; aliked ; glfeat .
Local Feature Matching: Before deep learning stage, researchers often apply the nearest neighbor (NN) searching to find matches between the extracted local features and perform ratio test to reject mismatches sift . This strategy is then improved by using the Earth Mover’s Distance Pele2008ECCV . Then, FLANN algorithm FLANN is developed as a efficient alternate of brute-force searching. To get more robust matches, various methods explores local key points distribution Guo2012PRL ; Ma2014TIP ; LPM ; Ma2018GLP ; CODE ; GMS . In addition to local key points distribution exploration, other strategies such as suitable local feature selection Hu2015TIP and optical flow guided matching Maier2016GuidedMB have also been explored to recover more matches. However, the above methods could ignore the global consistency since they only focus on the local distribution. On the other hand, the global properties of key points have also been explored ICF ; Yaron2014ToG ; RIRW ; DCM ; SOGM ; Lorenzo2008ECCV ; BBHomo . Moreover, the feature matching can be integrated into a RANSAC loop ransac to obtain matches with most inliers. Recently, AdaLAM adalam has been proposed to detects inliers by searching for significant local affine patterns in image correspondences. However, these traditional matching methods are less computational efficient due to hand-crafted rules and iterative formulation, and the performance will degrade if the underlying matching model between local feature in real scenarios differs from the pre-defined model of matching algorithms. In VPS , Sattler et al. proposed a novel vocabulary-based prioritized matching step that enables to first consider features more likely to yield 2D-to-3D matches and to terminate the correspondence search as soon as enough matches have been found, making VL-MRL method much more efficient and effective.
As for learning-based local feature matching methods, Some algorithms explore semantic information to match the semantic key points in different instances of the same category of objects DeepSemanticFM ; Khan2020TCSVT ; Yu2018HierarchicalSemanticIM . Inspired by PointNetPointNet , a DNN of the point set learning, putative matches filtering networks good_correspondence ; NMNet ; OANet ; ACNe are developed. They commonly organize the putative matches of key points as 4D quads and feed the 4D quads into network to estimate scores of each putative match. False matches are discarded by a threshold on the scores. These methods merely consider the position information of the key point and neglect the descriptor of local features.
In 2019, SuperGlue superglue was proposed as a E2E learning-based approach for local feature matching. SuperGlue superglue takes advantage of both key points and descriptors of local features and learns their matches with a graph neural network (GNN). The descriptor of each local feature is refined with all the features in both images by a attentional GNN layer so that matching will be reliable. Later, Zhao et al. IMUPrior exploited motion estimation from IMU integration and use the estimation as a spatial distribution prior of key points between images. With the assistance of such a prior, the network can make less efforts to refine features and achieve comparable accuracy with less GNN layers, boosting the real-time performance. More recently, LightGlue lindenberger2023lightglue improves the inference speed by enabling the model prune unmatchable points and predict whether further computation is required.
Summary: The local feature extraction-then-matching scheme has worked well in most cases and supported high-precision VL-MRL for a long time. However, there remains a fundamental shortcoming that they cannot match well if the extracted local features were extremely bad. The supervision on matches can only optimize matching module but cannot make any effort on the extraction module, as shown in Fig. 9 a). To solve this problem, joint local feature extraction-and-matching solution, or called detector-free matching, is proposed.
4.2 Joint Local Feature Extraction and Matching
Technically similar to SIFT Flow SIFTFLow , joint local feature extraction-and-matching methods integrate feature extraction and matching module in an unified DNN model, and they directly estimate dense local feature matches so that the two module can be jointly optimized, as shown in Fig. 9 b). Early proposals in this area adopt convolutional DNN (CNN) model, e.g., NCNet NCNet , Sparse NCNet SparseNCNet , and DRC-Net DRCNet . Although all the possible matches should be considered in the CNN of these works, the receptive field of convolution is still limited to neighborhood area, making matching sided and unliable.
Figure 9: The diagrams of a) local feature extraction-then-matching and b) extraction-and-matching methods.
To solve the problem, in recent three years, some works begin to adopt global attention mechanism, Transformer transformer , to achieve global consensus between matches. As the first proposal, COTR COTR applies zooming in method to obtain more accurate matching. LoFTR LoFTR proposes to perform dense matching in a coarse-to-fine framework, that it firstly obtains confidence matrices for patch-level coarse matching and then refines matched points positions at pixel level. Both steps using self- and cross-attention layers in Transformer transformer to enhance the representative ability of feature descriptors. It achieves accurate matching performance even in textureless region, as shown in Fig. 10, effectively reduce the requirements of scene texture in VL-MRL methods. As follow-up methods, QuadTree LoFTR Tang2022QuadTreeAF , ASpanFormer Chen2022ASpanFormerDI , and TKwinFormer Liao2023TKwinFormerTK make efforts to improve the matching accuracy and efficiency.
Figure 10: Some instances of local feature matches by SuperPoint superpoint +SuperGlue superglue , DRCNet DRCNet , and LoFTR LoFTR . The picture is from LoFTR .
Existing matching methods choose to ignore the occluded objects caused by camera motion and scene structure, and they will fail when there is no overlap region between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. To overcome this limitation, $\mbox{OCC}^{2}\mbox{Net}$ Fan2023Occ2NetRI models occlusion relations using 3D occupancy, enabling local feature matching for occluded regions.
Summary: The joint extraction-and-matching scheme breaks through the limits that matching need to first extract excellent local features, and achieves impressive matching performance even in challenging scenarios, making VL-MRL methods more practical. But it need both $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ as inputs, which requires the scene map additionally saves $\mathcal{I}_{r}$, enlarging the storage cost of scene map. A storage-friendly scheme need to be explored.
4.3 Pose Solver
Given the camera instrinsic $K$ and camera pose $\textbf{x}^{\mathcal{G}}_{q}$, the 2D camera projection $\textbf{p}^{\mathcal{I}_{q}}_{i}=\left[u_{i},v_{i}\right]$ from 3D visual landmarks $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},*>,\textbf{p}^{\mathcal{G}}_{i}=\left[X^{\mathcal{G}}_{i},Y^{\mathcal{G}}_{i},Z^{\mathcal{G}}_{i}\right]$ to image plane $\mathcal{I}_{q}$ can be formatted as:
$\begin{split}\left[\begin{array}[]{c}\textbf{p}^{\mathcal{I}_{q}}_{i}\\ 1\end{array}\right]&=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[X^{\mathcal{C}}_{i}~{}Y^{\mathcal{C}}_{i}~{}Z^{\mathcal{C}}_{i}\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[\textbf{R}^{\mathcal{C}}_{\mathcal{G}}~{}|~{}\textbf{t}^{\mathcal{C}}_{\mathcal{G}}\right]\left[X^{\mathcal{G}}_{i}~{}Y^{\mathcal{G}}_{i}~{}Z^{\mathcal{G}}_{i}~{}1\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\left[\begin{array}[]{c}\textbf{p}^{\mathcal{G}}_{i}\\ 1\end{array}\right]\end{split}$ (12)
where {}${\{\cdot\}}^{\mathcal{C}}$ denotes points in camera frame $\mathcal{C}$. And PnP problem is to estimate $\hat{\textbf{x}}^{\mathcal{G}}_{q}$ given $K$ and $n$ 2D-3D correspondences $\{\textbf{p}^{\mathcal{I}_{q}}_{i},\textbf{p}^{\mathcal{G}}_{i}\}^{n}_{i=1}$. This is a typical problem in 3D vision. Direct Linear Transformation (DLT) was first developed by photogrammetrists DLT to solve PnP problem and then introduced to computer vision community MVG ; 3DCV . If the 3D visual landmarks are coplanar, the homography transformation can be used instead MVG . In the last two decades there has been a great deal of progress in PnP solvers. The minimum number of correspondences to solve PnP prblem is reduced from 5 P5P , 4 P4P , to 3 p3p2 ; p3p1 . And the prominent solver is P3P algorithm ransac ; p3p2 ; p3p1 . Most of the solvers are able to deal with an arbitrary number of points. These generic PnP solvers are categorised as non-iterative and iterative.
Early non-iterative methods are generally computationally expensive, e.g., $O(n^{5})$ Quan1999TPAMI and $O(n^{2})$ Fiore2001TPAMI . The EPnP solution Lepetit2009EPnPAA is the first widely known $O(n)$ solver for the PnP problem. It represents 3D point coordinates as a linear combination of four control points. These control points are then retrieved using linearization techniques. Apart from linearization techniques, polynomial solvers have also been explored and become mainstream RPnP ; DLS ; OPnP ; SRPnP ; EOPnP . These works all estimate the camera pose by solving polynomial equations with $O(n)$ complexity.
Iterative methods are used to solve PnP optimization problems in an iterative manner, and it requires an appropriate initial values LHM ; FP ; SQPnP . The non-iterative solutions can be employed to provide an initial values, and iterative algorithms, such as Gauss-Newton (GN) iterations, are then used to refine the estimation. For instance, EPnP Lepetit2009EPnPAA and MLPnP Urban2016MLPnPA both enhance the performance by adding the GN iterations.
The PnP problem can also be seen as a non-linear least square problems with regard to re-projection error so that camera pose can be solved by Bundle Adjustment (BA) orbslam3 . Looking back to Eqn. 12, the re-projection error can be formatted as:
$\begin{split}e_{i}=\overline{\textbf{p}}^{\mathcal{I}_{q}}_{i}-\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\overline{\textbf{p}}^{\mathcal{G}}_{i}\end{split}$ (13)
where $\overline{\textbf{p}}$ means homogeneous coordinates of p. To estimate optimal camera pose, we need to minimize the total re-projection errors:
$\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{q}}\frac{1}{2}\sum^{n}_{i=1}e^{2}_{i}\end{split}$ (14)
And the function can be iteratively done by non-linear optimization methods like GN method and Levenberg-Marquardt (LM) method. In ORB-SLAM3 orbslam3 , the VL-MRL after tracking lost or kidnapping is implemented by first using MLPnP solver Urban2016MLPnPA to estimate an initial result and then refining with BA optimization.
Summary: Given accurate 2D-3D correspondences, PnP is a well-established problem and can be excellently solved. Pose can be precisely estimated. However, in most real-world cases, 2D-3D correspondences usually have many or even predominant mismatches, challenging the pose estimation procedure. The PnP solver is always integrated into a RANSAC loops ransac to estimate more robustly and precisely hloc2 ; orbslam3 .
4.4 Further Improvements
There are several other attempts to improve VL-MRL methods, except feature extraction and matching, and pose solver. Here we introduce some works with noticeable improvements.
Cross Descriptor matching cross-desc : Traditional VL-MRL methods supposes local features used in mapping and localization stage are same, which blocks the usage of new local feature algorithm. For example, if we used SIFT sift in mapping while HardNet hardnet in localization, traditional VL-MRL solution fails. In cross-desc , the authors overcame this problem by proposing to translate different type of descriptors into other descriptors or joint embeddings so that different descriptors can be indirectly matched.
Line Feature PL-Loc : Along with point-wise local features for matching, line-wise local features provide additional constraints in VL-MRL methods, especially in human-made environment. In PL-Loc , authors first presented a novel line segment descriptor and suggested the line feature can be complementary constraints to traditional point feature, as show in Fig. 11, thereby boosting localization performance on extreme scenarios where point features are biased or sparse.
Figure 11: Points and lines are complementary for better localization PL-Loc .
Dense CNN Matching InLoc ; PoseVerification ; Stumberg2019GNNetTG : Indoor environments are full of large textureless areas, where sparse local feature extraction methods detect very few features and extracted local features are clustered in small, well-textured parts in the image, leading to potentially unstable configurations for VL-MRL. To overcome this problem, InLoc InLoc and its extension PoseVerification use multi-scale dense CNN features for both image description and feature matching. GN-Net Stumberg2019GNNetTG is optimized with the proposed GN loss for training weather-invariant deep features, which is tailored for direct image alignment.
VL-MRL without SfM meshloc : Traditional VL-MRL need to reconstruct scene map using SfM colmap in offline mapping stage, which is computational and memorial costly. MeshLoc meshloc explores a more flexible alternative based on dense 3D meshes that does not require local features matching between $\mathcal{I}_{r}$ to build the scene map. During localization stage, MeshLoc meshloc still perform local feature matching on 2D images and estimate pose by solving PnP prblem, but the only difference is $\mathcal{I}_{r}$ is online rendered by 3D mesh model.
Map Squeeze SceneSqueezer : Storing pre-built scene map with visual landmarks can be prohibitively expensive for large-scale environments, especially on mobile devices with limited storage and communication bandwidth. In SceneSqueezer , authors designed a novel framework, SceneSqueezer, to compresses a scene while still maintaining localization accuracy. The scene $\mathcal{M}$ is compressed in three stages: 1) $\mathcal{I}_{r}$ for mapping are clustered; 2) a learned point selection module prunes the visual landmarks in each cluster taking into account the final pose estimation accuracy; 3) the descriptors of the selected visual landmarks are further compressed using learned quantization. By applying SceneSqueezer SceneSqueezer , the size of scene map is reduced to 31 MB compared to 7828 MB of HLoc hloc2 in Aachen Day-Night dataset benchmark6dof while achieving comparable localization performance.
Pose Verification InLoc ; PoseVerification and Correction PoseCorrection : Some VL-MRL methods estimate a set of candidate poses for $\mathcal{I}_{q}$, a pose selection procedure need to be conducted. InLoc InLoc utilizes the 3D structure of indoor scene and explicitly renders a virtual view that shows how the scene looks like from the estimated query pose. Then, authors extracted RootSIFT RootSIFT on a regular grid or dense CNN features from synthesized view and measured the region similarity between $\mathcal{I}_{q}$ and synthesized image, which can be seen as a score of candidate poses. In PoseVerification , Taira et al. improved pose verification by integrating semantic constraint and trainable pose verification network. In PoseCorrection , to circumvent the sparsity of co-visible local features between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ and improve the accuracy, Hyeon et al. proposed pose correction to reorganize local features observed from the estimated pose so that more reliable candidate poses can be provided to pose selection.
Open-sourced Toolbox hloc2 ; Kapture ; xrlocalization : To facilitate the development of VL-MRL researches, many great works have open-sourced the overall localization pipeline for the community, e.g., HLoc [^2] https://github.com/cvg/Hierarchical-Localization, Kapture-Localization [^3] https://github.com/naver/kapture-localization, and XRLocalization [^4] https://github.com/openxrlab/xrlocalization. By leverage these toolbox, researchers can quickly valid their new contributions or evaluate their new proposals without any unrelated efforts to build localization pipeline.
5 Point Cloud Map
Visual landmark map in Sec. 4 has a large requirement of storage due to saved high-dimensional features descriptors. The sensitivity of appearance (e.g., illumination) also challenges mapping with monocular camera and affect the reconstruction of visual landmark map. As a alternate, LiDAR can provide direct and precise 3D perception of scene map geometry, and the point cloud map built by high-precision LiDAR always be illumination-invariant and commonly more accurate than visual map. And point cloud map only need to save 3D position of point cloud, alleviating the storage burden. The point cloud map can be represented as =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a 3D point in global frame $\mathcal{G}$, $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},\textbf{l}_{i}>$, $\textbf{p}^{\mathcal{G}}_{i}$ is its 3D coordinates and $\textbf{l}_{i}$ is its intensity (intensity sometimes is missing). In practical usage, the 3D points are usually used in “clip” format, that is, given a virtual viewpoint $\textbf{x}^{\mathcal{G}}_{v}$, the 3D point observed in this view can be clustered in a set $\mathcal{P}^{\mathcal{G}}_{v}=\{\textbf{p}_{i},\textbf{l}_{i}\}$, so $\mathcal{M}$ is changed to =$\mathcal{M}=\{\textbf{m}_{i}\}$ where $\textbf{m}_{i}$ is a clip of 3D point and its associated viewpoint $\textbf{m}_{i}=<\mathcal{P}^{\mathcal{G}}_{v},\textbf{x}^{\mathcal{G}}_{v}>$. Since the wide deployment of LiDAR localization is hampered by the huge cost of high-precision LiDAR, localizing camera in LiDAR point cloud map (named as Image-to-Point cloud (I2P) localization for simplicity) is a newly emerged trend in MRL. However, the appearance information of camera and the geometry information of LiDAR is in different modality, the inherit difficulties in cross-modal data association challenges I2P Localization. In this section, we review cross-modal I2P localization using traditional geometry rules and using learning-based methods.
Figure 12: The diagram of four typical cross-modal I2P localization methods. including geometry-based a) G-I2P mono_i2p_vo , and learning-based b) I2P-VPR, c) I2P-RPR, and d) I2P-MLoc.
5.1 Geometry-based Cross-modal Localization
Geometry-based I2P (G-I2P) methods localize camera in LiDAR map by traditional geometry rules. In the G-I2P researches, Stereo camera is a natural pick because we can easily lift 2D image data to 3D by stereo matching, which makes the alignment between camera and point cloud map much easier stereo_i2p . Instead, Localizing monocular camera images in 3D point cloud maps is inherently more challenging, as they lack any depth or 3D information. For this, researchers turn to align $\mathcal{I}_{q}$ to $\mathcal{M}$ by lifting 2D visual measurements to 3D local construction mono_i2p_vo , digging geometry features mono_i2p_line ; Wolcott2014IROS , or modelling 3D structure in other way mono_i2p_SDF ; dsl ; Huang2020GMMLocSC . As an early and successful attempt, Caselitz et al. mono_i2p_vo proposed to associate the visual landmarks reconstructed from the monocular visual odometry orbslam with point cloud map, as show in Fig. 12 a). The 7 DoF transform was estimated in an ICP scheme. In Wolcott2014IROS , authors proposed to render synthetic images using intensity acquired from the LiDAR sensor so that pose estimation can be performed by maximizing the normalized mutual information between real $\mathcal{I}_{q}$ and synthetic images. Although identifying matched features across different modalities is proven challenging, the incorporation of line features has been demonstrated to aid in this objective in mono_i2p_line . Once the authors identified the co-visible lines in both 3D and 2D views, they designed a feature vector for each line, which incorporated various properties like length and orientation. Then, line feature matching between 2D and 3D views is performed to provide geometric constraint for pose estimation. Huang et al. modelled the dense structure as an Euclidean Signed Distance Field (ESDF) mono_i2p_SDF . The visual structure can then be aligned with the implicit surface. Ye et al. proposed DSLdsl where surfel constraints of point clouds are introduced into the direct photometric error. The monocular camera can be localized in a tightly-coupled BA framework with photometric error. More recently, GMMLoc Huang2020GMMLocSC uses a Gaussian Mixture Model (GMM) GMM to model the prior distribution of the point cloud map. Assisted by the GMM map, the performance of baseline orbslam2 in localization is significantly enhanced in challenging situations.
Summary: Many attempts on G-I2P have been made and achieved impressive successes, but all of these works need to be integrated into a VO pipeline to obtain initial guess of pose, and they are hard to re-localize with only one monocular image. These methods tend to apply continuous re-localization in point cloud map for absolute pose estimation in order to add absolute constraint into VO system, alleviating the drift problem of long term exploration.
5.2 Learning-based Cross-modal Localization
Cross-modal Visual Place Recognition (I2P-VPR): Recently, scholars have begun to explore deep learning-based methods to achieve I2P localization with only one monocular image. First, many researches adopt GF-VPR method (in Sec. 3.1) into this cross-modal I2P localization task and propose I2P-VPR algorithms. I2P-VPR methods are very similar to GF-VPR methods, but its reference “frames” $\mathcal{I}_{r(t)}$ is changed to local clips of point cloud map $\mathcal{P}^{\mathcal{G}}_{r(t)}$. For clearness, we also provide detailed pipeline of I2P-VPR methods here. Given the query image $\mathcal{I}_{q}$, a I2P-VPR algorithm can be written as a two-step pipeline, that is,
  • 1
    represent $\mathcal{I}_{q}$ and all the reference point cloud clips $\mathcal{P}^{\mathcal{G}}_{r(t)}$ in $\mathcal{M}$ as features:
[…]
  • 2
    retrieve the best-matched point cloud clip based on the similarity of features:
    $\hat{\textbf{m}}=\mathop{\arg\max}\limits_{\mathbf{m}_{i}}P(\{\mathcal{F}^{q},\mathcal{F}^{i}~{}\mbox{of}~{}\textbf{m}_{i}|\textbf{m}_{i}\in\mathcal{M})$ (16)
Then, the pose of $\mathcal{I}_{q}$ can be approximated as the associated pose with $\hat{\textbf{m}}$.
For this problem, one straightforward solution is jointly training a 2D CNN $F_{\mbox{ImFeat}}(\cdot)$ for images and a 3D DNN $F_{\mbox{PcFeat}}(\cdot)$ for point clouds to create shared embeddings 2d3d_embed . However, this approach does not generalize well to unseen environments. Lukas2021ICRA_Spherical proposes to project images and point clouds into unit spheres and extract features through sphere CNN, which requires multiple images as input. Towards I2P localization robust to inconsistent environmental conditions, i3dLoc i3dloc matches equirectangular images to the 3D range projections by extracting cross-domain symmetric place descriptors. The 3D geometry features eliminates the condition-related features by a designed Generative Adversarial Network. Based on such features, i3dLoc i3dloc further design a spherical convolution network to learn viewpoint-invariant symmetric place descriptors. Similarly, AE-CrossModal focuses on correlating the information of 360-degree spherical images to point clouds. Attention mechanism is applied to let the network capture the salient feature for comparing images and point clouds. I2P-Rec Li2023I2PRecRI provides a new baseline for I2P-VPR that it leverages on depth estimation networks to recover point clouds from images so that cross-modal data is converted into the same modality, i.e., Bird-Eye-View (BEV) images. Using the BEV image as an intermediate representation, very simple global feature network (CNN encoder followd by NetVLAD layer) trained by small number of training data can achieve impressive localization performance.
Cross-modal Relative Pose Regression (I2P-RPR): After obtaining matched point cloud clip $\mathcal{P}^{\mathcal{G}}_{r}~{}\mbox{of}~{}\textbf{m}_{i}$ and its pose $\textbf{x}^{\mathcal{G}}_{r}$, many works further refine query pose by I2P-RPR that is similar to image-to-image RPR in Sec. 3.2, that is, estimate relative pose between $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$:
$\hat{\textbf{x}}^{r}_{q}=\mbox{RPR}(\mathcal{I}_{q},\mathcal{P}^{\mathcal{G}}_{r})$ (17)
so that the absolute pose of $\mathcal{I}_{q}$ can be calculated by:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}\leftarrow\textbf{x}^{\mathcal{G}}_{r}*\hat{\textbf{x}}^{r}_{q}$ (18)
[…]
However, I2P-RPR methods are hard to achieve centimeter-level accuracy on large scale scene. So many proposals turn to re-localize camera based on 2D-3D matching between images and point cloud map, which is technically similar to VL-MRL methods in Sec. 4 and is also called image-to-point cloud registration.
Cross-modal Matching-based Localization (I2P-MLoc): 2D3D-MatchNet 2d3dmatchnet is one of the earliest works focusing on image-to-point cloud registration for robot localization. The 2D and 3D key points are obtained by SIFT sift and ISS ISS respectively. Then, a neural network is introduced to learn the descriptors for key points. Finally, EPnP Lepetit2009EPnPAA is adopted to estimate the transformation between the image and the point cloud with the 2D-3D correspondences. Built upon CMRNet cmrnet , CMRNet++ cmrnet++ turn to regress displacements between images and projected depth of point cloud map instead of directly regressing pose. CMRNet++ cmrnet++ achieves improved performance with significant margin, indicating that matching-based methods should be a more reasonable choice than regression-based alternates for I2P localization. DeepI2P deepi2p splits the image-to-point cloud registration problem into a classification problem and an optimization problem. A cross-modality neural network is adopted to classify whether the points fall into the image frustum. The classification results are utilized to construct the cost function of the inverse camera projection. The optimization solver solves the transformation that minimizes the value of the cost function. CorrI2P CorrI2P designs a cross-modality network to extract the image-to-point cloud overlapping region and corresponding dense descriptors for the image and point cloud. CorrI2P CorrI2P constructs dense image-to-point cloud correspondences and uses iterative RANSAC-based EPnP Lepetit2009EPnPAA ; ransac to estimate the relative pose. EFGHNet EFGHNet adopts the divide-and-conquer strategy to divide the image-to-point cloud registration into four separate sub-networks. These sub-networks are responsible for the horizon and ground normal alignments, rotation estimation, and translation estimation, respectively. These mentioned above methods divide the image-to-point cloud registration into separate modules for robot localization. The separation makes the modules separately optimized and thus not able to refine the error of the previous modules. Thus, I2D-Loc I2DLoc apply BPnP module BPnP to calculate the gradients of the back-end PnP-based pose estimation process, enabling the model to be trained end-to-end with the supervision on estimated pose. More recently, I2PNet I2PNet utilizes a coarse-to-fine 2D-3D registration architecture to accurately regress pose. By combining both advantages of regression-based and matching based I2P localization methods, I2PNet I2PNet achieves centimeter-level localization accuracy.
Some other works add visual features in point cloud map, which is quite similar to visual landmark map in Sec. 4 but the visual features of point cloud in map are online extracted during localization without offline storage. NRE introduces the Neural Re-projection Error (NRE) as a substitute for re-projection error in PnP problem. Given $\mathcal{I}_{q}$ and $\mathcal{P}^{\mathcal{G}}_{r}$ in map with $\mathcal{I}_{r}$, the method extracts dense CNN features from both images. The sparse descriptors of 3D point clouds are sampled from $\mathcal{I}_{r}$ based on image projection as Eqn. 12. For each 3D point, NRE computes dense loss maps and minimizes the loss with respect to the network parameters. The proposed NRE in NRE is differentiable not only w.r.t. the camera pose but also w.r.t. the descriptors. PixLoc pixloc is another work in this field. PixLoc pixloc also extracts multi-level features with pixel-wise confidences for $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. The LM optimization then aligns corresponding features according to the 3D points in the map, guided by the confidence, from the coarse to the fine level. Compared with GN-Net Stumberg2019GNNetTG that also trained deep features via direct alignment, PixLoc pixloc leverages the power of differentiable programming, and the only used supervision in training is pose, without the requirement of accurate pixel-wise ground truth correspondences.
Summary: The inherit difference between cross-modal data make I2P localization hard. Benefit from high-dimensional CNN features, matching across cross-modal data becomes possible. In general, I2P-VPR only serves to retrieve reference point cloud clip and provides an coarse pose estimation. I2P-RPR and I2P-MLoc could provide more precise pose estimation. I2P-RPR avoid explicit feature matching and pose solver, the training costs is cheap but it cannot achieve very high localization accuracy. Instead, I2P-MLoc performs best in cross-modal localization task, showing good prospects for development.
6 Vectorized High-Definition Map
In AD task, the compactness and accuracy are two vital characteristics of scene map. Therefore, the most widely used map in AD area has been HD Map since the its first appearance in late 2010. Commonly, the HD Map is created by mobile mapping system equipped with high-precision sensors including LiDAR, RTK and IMU at centimeter-level precision HDMap_review . The localization feature in the HD map includes dense point cloud and sparse map element. The dense point cloud in HD Map is similar to the raw point cloud map in Sec. 5. In this section, we focus on the sparse map elements usually represented as vectors with semantic labels. These elements correspond to road element or signatures in the real world, e.g., lighting poles, road markers, and lane lines. Such a semantic element-based map representation is much lighter than other scene map like point cloud map and visual landmark map while carries road elements with a high level of detail HDMap_review . MRL with HD map (HD-MRL) is concluded as an effective solution for mass-produced vehicles. The fundamental formulation of HD-MRL methods involve identifying semantic elements of HD map from images and then estimating pose by aligning the detected 2D element in the image with their corresponding 3D element in the HD map.
Figure 13: A diagram of MRL method with vectorized HD map.
[…]
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{q}}\sum_{\textbf{m}_{i}\in\mathcal{M}}\sum^{N_{i}}_{j=1}d\left(\textbf{z}^{\mathcal{I}_{q}}_{i(j)},\tilde{\textbf{p}}^{\mathcal{I}_{q}}_{i(j)}\right)$ (19)
where $\textbf{z}^{\mathcal{I}_{q}}_{i(j)}$ is the observation of $j$-th point of $i$-th element $\textbf{m}_{i}$ in image plane $\mathcal{I}_{q}$.
From the basic formulation defined above, HD-MRL methods need to be integrated into a localization system with multiple sensor fusion for acquiring an initial camera pose, which is similar to G-I2P in Sec. 5.1, since HD map elements can be seen as sets of 3D point cloud with semantic information. In this field, Pink et al. Oliver2008CVPRW used aerial images to extract lane markings and their respective positions to build a lane-level map, then matched the lane marking in the image with this pre-built map using ICP algorithm. The subsequent works improve HD-MRL by fusing additional sensors such as GNSS and IMU in an Extended Kalman Filter (EKF) framework. Tao et al. Tao2013IROS proposed to use the EKF fusion framework to integrate lost GPS, dead reckoning, and visual features from lane markings. The lane markings detected by the video camera system provide the lateral distance and heading angle information of a vehicle. Cai et al. Cai2018Sensors also used the matching between 2D visual observation and 3D HD map element as the additional lateral distance observation in the filter-based fusion frameworks.
Figure 14: Examples of alignment between HD map and image semantic segmentation Guo2021IROS . (a) An accurate alignment; (b)-(g) Alignment results under perturbation of freedom: roll, pitch, yaw, x, y and z. Semantic segmentation: freespace (yellow), lane markings (organ and purple). Projection from HD map: lane markings (green) and signboard (cyan).
[…]
Another main problem blocking the usage of HD-MRL methods is the the sparsity of HD map landmark features. To solve this problem, multiple sensor fusion often be used in EKF frameworks that IMU and GPS are utilized for continuous localization Tao2013IROS . With the development of VO vins-mono and visual(-inertial) SLAM orbslam ; orbslam2 ; orbslam3 , visual features are shown to be able to serve as reliable constraints for relative pose estimation. They can be considered as complementary localization information in conjunction with the HD map landmark features. Some solutions combine these two pieces of information in a loosely coupled manner, i.e., the HD MRL and the visual feature-based localization results are calculated separately first, and then fused by Kalman Filter Julius2014IV or sliding window-based state estimator MLVHM . However, these loosely-coupled strategies share a common shortcoming: they all require the HD map elements to be minimally self-sufficient for solving vehicle poses, which cannot be satisfied in real-world applications. Wen et al. Wen2020IV ; TM3Loc followed this way to combine VO and HD-MRL but in a tightly-coupled framework with sliding window-based optimization. The basic concept is to optimize the poses with the minimization target consisting of both visual feature residuals from VO and HD map element residuals from HD-MRL, enabling consistent and accurate localization even when HD map elements are insufficient in current observation.
In last two years, some deep learning-based HD-MRL methods have been developed. OrienterNet sarlin2023orienternet estimates 3 DOF pose in a 2D map, OpenStreetMap (OSM), by proposed neural map matching. But its localization accuracy is only limited to sub-meter level due to the low precision of used OSM map. BEV-Locator bevlocator formulates this problem as an E2E learning scheme and proposes a Transformer-based architecture to address the key challenge of the cross-modality matching for HD-MRL. It encodes the map elements from discrete points into structured vectors, and conduct interaction between images and HD map on BEV space. EgoVM he2023egovm extracts BEV features and map embeddings by Transformer Decoder transformer , and then map embeddings are compared with interpolated map features by candidate poses to calculate their similarities, so as to estimate the optimal pose offset. However, BEV-Locator bevlocator and EgoVM he2023egovm needs multiple view images as input, learning-based HD-MRL problem still be unsolved.
Summary: HD map should be the most compact and lightwight scene map, so it is preferred by AD researches. Although it need heavy construction and maintenance cost, the recently proposed crowdsourced map building crowdsourcedmapping and updating GCDL strategies may help to reduce the maintenance cost, making HD map more widely acceptable. MRL in HD map is still a very hard problem because precise data association between visual observations and semantic elements in HD map is theoretically difficult. Traditional methods seek aids from multiple sensor fusion or online odometry while recent learning-based methods leverage advanced representation ability of DNN models. Based on the localization performance and required sensor setup, we believe learning-based method is a more promising way for HD-MRL.
Figure 15: Diagrams of MRL method with implicit map, such as a) APR methods, b) SCR methods, and c) and d) NeRF-based methods.
7 Learnt Implicit Map
In the current age of deep learning, scholars begin to rethink the representation format of scene map in a more implicit way. Some recently proposed works encode scene map into neural networks so that the network can achieve amazing things, such as directly recover pose of images (called Absolute Pose Regression, APR), estimate 3D coordinates of each pixel in images (called Scene Coordinate Regression, SCR), or render geometry structure and appearance of scene (called Neural Radiance Field, NeRF). Benefit from high-performance computation device, matured deep learning technology, and tremendous number of training samples, such implicit map may be new choice for MRL methods (called NN-MRL) in the near future. In this section, we review APR, SCR, and NeRF-based MRL methods, as shown in Fig. 15.
7.1 Absolute Pose Regression
APR solution was first proposed by PoseNet posenet to directly regress the position and orientation of the camera with regard to global frame, given a query image $\mathcal{I}_{q}$, by attaching an MLP head to a GoogLeNet backbone googlenet . Although early APR methods achieve less accuracy than the VL-MRL methods (Sec. 4), but enable pose estimation using a single forward pass. The scene map is represented in a totally implicit and very compact formulation, i.e., parameters of DNN $F_{\mbox{APR}}(\cdot)$. Thus, APR method can be simply formatted as:
$\hat{\textbf{x}}^{\mathcal{G}}_{q}=F_{\mbox{APR}}(\mathcal{I}_{q})$ (20)
Single Scene APR (SS-APR): In order to improve the localization accuracy, contemporary APR methods study different designs of CNN backbones Shavit2020CoRR ; Tayyab2017IROS ; hourglassPN , MLP heads Tayyab2017IROS . These models can work very well on training scene but achieve limited accuracy on test scene. Such an overfitting problem is addressed by averaging the predictions of multiple models with randomly dropped activations BayesianPN , or by reducing the dimensional of the global image encoding using Long-Short-Term-Memory (LSTM) layers LSTMPN . Other work focused on the supervision strategy for absolute pose regression to allow adaptive weighting of position- and orientation-associated errors. In original PoseNet posenet , the scale between orientation loss and translation error varies across scenes, e.g., 120 to 750 for indoor scenes while 250 to 2000 for outdoor scenes, which is tricky. Kendall et al. PoseNet2 suggested optimizing the parameters that balance both losses to improve accuracy and avoid manual fine-tuning. This formulation is widely adopted by later APR methods. Alternative orientation representations were also proposed to improve pose loss mapnet . The use of geometric constraints from additional sensors, such as VO and GPS, is also suggested to be formulated as loss terms in network training to improve localization accuracy mapnet . In Saha2018ImprovedVR , SS-APR problem is split into two step that first predicts the most relevant anchor point uniformly defined in the scene and then regress the relative offsets with respect to it. Such a scheme simplifies the APR problem and significantly improve the accuracy. More recently, Wang et al. Wang2019AtLocAG proposed using attention to guide the pose regression process by applying self-attention to the CNN features. CaTiLoc catiloc exploits attention mechanism on full receptive field by Transformer transformer . Xue et al. LSG incorporated learning-based VO estimation and APR in a single DNN and used joint training for RPR and APR so that spatial-temporal consistency can be leveraged to alleviating the uncertainty of pose estimation by augmenting the observation based on the co-visibility from VO estimation. In testing stage, Brahmbhatt et al. mapnet and Xue et al. LSG performed a standard Pose Graph Optimization (PGO) by optimizing only the nodes (APR estimation) with edges (VO estimation) fixed.
Although many modifications are proposed to the architecture and training method originally formulated by Kendall et al. posenet , the main paradigm remained the same that employs (1) a CNN backbone to generate a global latent feature, and (2) a regressor to estimate absolute pose based on extracted features. The backbone and regressor are E2E trained for each scene in the SS-APR methods mentioned above.
Multiple Scene APR (MS-APR): MS-APR methods are later developed aiming to extend a single trained APR model on multiple scenes. Blanton et al. MSPN first proposed Multi-scene PoseNet to first classifies the particular scene related to the query image, and then uses it to index scene-specific regressor to estimate the pose. They train a set of regressors, one per scene, for APR with a set of scene-specific parameterized losses. MS-Transformer MS-Transformer1 ; MS-Transformer2 then learns multi-scene APR with Transformers, where encoders are used to aggregate features with self-attention and decoders transform latent features and scenes encoding into pose predictions. MS-Transformer MS-Transformer1 ; MS-Transformer2 allows the model to focus on informative features for localization by stacked attention, while embedding multiple scenes in parallel.
Summary: Even APR algorithms work well on small scale dataset, e.g., 7Scenes 7scenes , there is no evidence that they can scale to large scale scene. For example, on Cambridge Landmarks dataset posenet that consists of six medium-sized scenes (about 900-5500 $\mbox{m}^{2}$) in an urban environment, the state-of-the-art (SOTA) MS-Transformer MS-Transformer2 achieves an average localization accuracy of 0.98 m and ${}^{\circ}$, much worse than VL-MRL method, InLoc InLoc , with accuracy of 0.11 m and ${}^{\circ}$. APR methods can serve as a initial or coarse pose estimation in some real-time application, and its $O(1)$ computational complexity is also remarkable.
7.2 Scene Coordinate Regression
SCR methods follow the PnP-based solution of MRL methods but in a E2E learning-based way. The scene map $\mathcal{M}$ is also represented as network parameters of DNN $F_{\mbox{SCR}}(\cdot)$. Specifically, SCR methods directly learn to predict 3D scene coordinates $\textbf{p}^{\mathcal{G}}_{i}$ of each 2D pixel $\textbf{p}^{\mathcal{I}_{q}}_{i}$ in the $\mathcal{I}_{q}$ and thus query pose can be solved as a PnP problem based on estimated 2D-3D correspondences, which can be formatted as:
$\begin{split}\{\hat{\textbf{p}}^{\mathcal{G}}_{i}\}&=F_{\mbox{SCR}}(\mathcal{I}_{q})\\ \hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{solvePnP}(\{\textbf{p}^{\mathcal{I}_{q}}_{i},\hat{\textbf{p}}^{\mathcal{G}}_{i}\}_{i})\end{split}$ (21)
where $\{\textbf{p}^{\mathcal{I}_{q}}_{i},\hat{\textbf{p}}^{\mathcal{G}}_{i}\}_{i}$ is a pair of image pixel and estimated scene coordinate.
In this field, early approaches SCRForests ; Abner2014CVPR ; Julien2015CVPR used regression forest models to predict 3D scene coordinates for images, but they need RGB-D images as input. More recent works DSAC ; DSAC2 ; HSCNet ; few-shot-SRC ; DSM ; KFNet ; SANet ; Less_is_More ; Huang2021VSNetVW have replaced regression forests with DNN-based models that only require a RGB image.
Scene-specific SCR (SS-SCR): SCR method is firstly used to estimate dense 3D coordinates for $\mathcal{I}_{q}$ in a previously seen scene, and SCR model need to be specifically trained for each scene, called scene-specific methods. Brachmann et al. DSAC ; DSAC2 ; ESAC ; Less_is_More trained neural networks for this task and combine them with a differentiable RANSAC strategy for MRL. For instance, in ESAC , multiple scene coordinate regression networks are trained as a mixture of experts along with a gating network which assesses the relevance of each expert for a given input, and the final pose estimate is obtained using a novel RANSAC framework, i.e., Expert Sample Consensus (ESAC). Dong et al. few-shot-SRC and Li et al. HSCNet later introduced region classification into their pipelines for effective scene memorization. Huang et al. Huang2021VSNetVW proposed to add a segmentation branch to obtain segmentation on scene-specific landmarks, which can then be associated with 3D landmarks in the scene to estimate camera pose. In Budvytis2019LargeSJ , SS-SCR is formulated as two separate tasks of object instance recognition and local coordinate regression. Since dense 3D structure of scene is required as ground-truth in model training, HSCNet++ Wang2023HSCNetHS removes the dependency on dense ground truth scene coordinates by proposing a pseudo-labelling method, where ground-truth labels at each pixel location are propagated to a fixed spatial neighbourhood. However, SS-SCR need to train models per scene and it commonly pixel-wise ground-truth scene coordinates, the heavy training costs block the wide usage of SS-SCR.
Scene-agnostic SCR (SA-SCR): To boost the generalization ability and scalability of SCR methods, SA-SCR methods are proposed SANet ; DSM ; Revaud2023SACRegSC . They can regress dense scene coordinates based on some given reference views in unseen scenes. SANet SANet first constructs a scene pyramid by extracting multiple scale feature map from reference images, then predicts scene coordinates for the query image with the assistance of scene pyramid. In DSM , authors proposed a Dense Scene Matching (DSM) module. The DSM module predicts scene coordinates by cost volume constructed between query image and reference image with its scene coordinates. Following this way, the Transformer-based SAReg Revaud2023SACRegSC takes a variable number of images (including query images and retrieved reference images) and sparse 2D-3D annotations from SfM database colmap as input.
More implicitly, NeuMap neumap encodes the scene as sparse scene-specific latent codes and estimates scene coordinates with the help of the resulting latent codes. When localizing in a new scene, NeuMap neumap only need to optimize latent code for the new scene without re-training other part of DNN. With the latent code pruning, redundant codes can be removed. So NeuMap neumap can achieves similar performance as DSAC++Less_is_More with a 100 to 1000 times smaller scene map size. Instead of densely estimate scene coordinates, NeuMap neumap and D2S Bui2023D2SRL estimate scene coordinates for sparse 2D key points, making them cost-effectiveness. D2S Bui2023D2SRL also supports updating by self-supervised learning procedure with new observations without requiring camera poses, intrinsic parameters, and ground-truth scene coordinates, which greatly improves the scalability and generalization ability of SCR.
Summary: SCR methods do not explicitly rely on local feature detection, description, and matching, and are able to provide dense correspondences between query image and scene map. They also do not require storing the 3D models and descriptors, as scene map is densely encoded in a learnable DNN. The localization accuracy of SCR is impressively great in small- and medium-scale scenarios 7scenes ; posenet , but its usage in unbounded scene is hard due to expensive training costs.
7.3 Neural Radiance Field
NeRF implicitly represents a scene as a “radiance field”: a volumetric density that models the geometry of the scene, and a view-dependent color that models the appearance of occupied regions of the scene NeRF . A Multiple Layer Perception (MLP) is utilized to takes the 3D position of a point $\textbf{p}^{\mathcal{G}}=(x,y,z)$ and the unit-norm viewing direction of that point $\textbf{v}=(d_{x},d_{y},d_{z})$ as input, and estimate the density $\sigma$ and color $c$ of the point, $(\sigma,c)\leftarrow F_{\mbox{NeRF}}(\textbf{p}^{\mathcal{G}},\textbf{v})$. Scene map is represented by the parameters $\Theta$ of $F_{\mbox{NeRF}}()$. Given a set of RGB images $\{\mathcal{I}_{t}\}$ with known camera poses $\{\textbf{x}^{\mathcal{G}}_{t}\}$, NeRF is trained by minimizing a photometric loss $\mathcal{L}$ NeRF :
$\hat{\Theta}=\mathop{\arg\min}\limits_{\Theta}\mathcal{L}(\{\mathcal{I}_{t}\},\{\textbf{x}_{t}\})$ (22)
NeRF learns to synthesize novel views associated with any camera poses, enabling virtual image rendering and scene geometry recovery, thus have been applied in MRL in recent years. Some works directly uses pre-trained NeRF to estimate pose while other works utilize NeRF as a novel data augmentation method to enrich the training data.
NeRF as pose estimator: As the first framework to estimate pose using NeRF, iNeRF [28] directly optimizes 6 DoF poses with freezed network parameters $\Theta$ by minimizing the photometric error between rendered pixels and observed pixels, which is the inverse problem of original NeRF NeRF :
$\hat{\textbf{x}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}_{q}\in SE(3)}\mathcal{L}(\mathcal{I}_{q},\textbf{x}_{q}~{}|~{}\Theta~{}\mbox{freezed})$ (23)
Such a joint pose estimation and NeRF training scheme then motivates pose-free NeRF Meng2021GNeRFGN ; NoPe-NeRF . In LATITUDE LATITUDE , authors trained a pose regressor through images generated from pre-trained NeRFs, which provides an initial value for hierarchical localization as VPR methods without the requirement of saving reference images. In fine pose optimization stage, LATITUDE LATITUDE minimizes the photometric error between the observed image and rendered image by optimizing the pose on the tangent plane. In Loc-NeRF LocNeRF , a pre-trained NeRF model is integrated into a Monte Carlo localization to assigns weights to particles by image similarity between query image and rendered image. In NeRF-Loc liu2023nerfloc , a NeRF conditioned on reference images takes 3D point as input and generates corresponding 3D descriptors so that 2D-3D matches can be obtained by matching. Pose is then estimated by PnP solver.
NeRF as data augmentation: As a more straightforward way to apply NeRF in MRL, NeRF can be used to render virtual images at unseen viewpoints so that the overfitting problem of APR methods can be effectively solved and thus improves performances of APR Moreau2021LENSLE ; DPN ; DFNet . LENS Moreau2021LENSLE applies pre-trained NeRF-W NeRF-W as an offline data augmentation pipeline to enhance APR model. As an online augmentation way, Direct-PoseNet DPN minimizes the photometric errors between query image and virtual image rendered at estimated pose while DFNet DFNet uses feature map errors DFNet during online localization at unseen viewpoints, making APR method more precise and scalable.
Summary: As a newly emerged technique, NeRF has two favourable characterises: 1) the implicit representation of scene; 2) the ability to render synthesis image at novel viewpoint. These two characterises present great potential in MRL researches. Current attempts in this area focus on incorporating NeRF with APR methods. With the assistance of NeRF, APR could perform comparable or even better than VL-MRL methods in small- and medium-scale scenes. But we have to say the utilization of NeRF in MRL is still an opening question to be explored.
Figure 16: Instances of APR method enhanced by NeRF, a) Direct-PoseNet DPN using photometric consistency between query image and rendered image, b) DFNet DFNet using feature consistency between query image and rendered image.
8 Benchmarking Monocular Re-Localization
In this section, the common evaluation metric and dataset are introduced, and then we provide the performances of some SOTA MRL methods with various type of scene map on these benchmark. Looking back to all kind of MRL methods mentioned above, some of MRL methods retrieve reference image as query pose approximation while others perform precise pose estimation for query image, as listed in Tab. 1. These two kinds of MRL are generally tested using different evaluation metric. We provide introduction and evaluation respectively.
8.1 Evaluation metrics
8.1.1 MRL  for  IPA
Those retrieval-based MRL utilize evaluation metric as image retrieval task. A retrieved reference image/point cloud clip is defined as true positive if it was geometrically near the query images. The ground-truth is generally labeled by geometric distance based on GPS signals (for example, a threshold of 10 m in Li2023I2PRecRI ) or manually defined yue1 ; yue2 . With the ground-truth, we can measure Precision and Recall metric, where precision denotes the ratio between true positives and all the retrieved references and recall is the ratio between the number of true positives and all the potentially matched query-references existed in this scene. In some MRL application need to retrieve comprehensively, we measure the ratio of retrieved true positives over top-K candidates, i.e., Recall@K. In some applications that a mismatch will make the whole system dramatically fails, e.g., loop closure detection in SLAM LCD2022TITS , we turn to measure the maximum recall metric with 100% precision. The higher these two metrics, the better MRL performs.
8.1.2 MRL  for  FPE
Direct measurement about estimated poses is applied in those MRL methods aiming at precise pose estimation. The common evaluation metrics are position error for translation $\textbf{t}^{\mathcal{G}}_{v}$ (in distance) and orientation error for rotation $\textbf{R}^{\mathcal{G}}_{v}$ (in degree) between estimated results and ground-truth poses. The lower the errors, the better MRL performs.
Additionally, the size of represented scene map $\mathcal{M}$ and the running efficiency of MRL methods should also be considered.
Table 1: The categorisation of various MRL methods based on the function.
Function
Method
Initial Pose Approximation
VPR (Sec. 3.1) and I2P-VPR (Sec. 5.2)
Fine Pose Estimation
RPR (Sec. 3.2), VL-MRL (Sec. 4), G-I2P (Sec. 5.1), I2P-RPR (Sec. 5.2), I2P-MLoc (Sec. 5.2), MRL with HD map (Sec. 6), APR (Sec. 7.1), SCR (Sec. 7.2), and NeRF-based MRL (Sec. 7.3)
8.2 Dataset
There are a large number of datasets available for the evaluation of MRL methods and it is difficult to give a complete overview. Here we only present some typical datasets that we have used to evaluate MRL methods in this survey.
8.2.1 MRL  for  VPR  only
The dataset used to evaluate VPR is required to include query and reference images that can be paired by geometric distance. So the dataset should have traverse the same scene multiple times, or the trajectory of the dataset should be “looped”.
Gardens Point dataset GardenPoint2 has been recorded on the Gardens Point Campus in Queensland University of Technology at different times of day and night. It contains three traverses of scene that two in the day while one in the night.
SPEDTest dataset SPED includes frames depicting various locations from all over the world. The collected scenes are with a vast range of weather, seasonal, and illumination conditions.
Nordland dataset nordland was first introduced to VPR evaluation by seqslam . It represents a 728 km of train journey in Norway during summer and winter seasons. Nordland dataset represents natural (non-urban), outdoor environment, which is unexplored in any other dataset.
Pittsburgh dataset pittsburgh has two versions of test dataset for VPR evaluation. Pitts250k-test contains 8k queries and 83k reference images, collected from Google Street View and Pitts30k-test comprises 8k queries and 8k references. Both Pittsburgh datasets show significant viewpoint changes.
Mapillary Street-Level Sequences (MSLS) dataset MSLS has been collected using camera mounted on ground vehicles and presents a wide range of viewpoint and illumination changes.
Oxford dataset fabmap , including CityCentre and NewCollege, is collected through the fish-eye binocular images of a wheeled robot in a campus. Ground-truth matches between query and reference image are provided by the authors.
St. Lucia dataset RatSLAM is recorded through a suburb in Brisbane at 5 different times during a day, and also on different days over a period of two weeks. The viewpoint changes are limited, but significant appearance changes due to the different times of day are included, as well as some dynamic objects such as traffic or cars parked on the street.
17-Places dataset 17places consists of a number of different indoor scenes, ranging from office to labs, hallways, seminar rooms, bedrooms, etc. This dataset exhibits both viewpoint- and conditional-variations.
Except to these datasets, tremendous challenging datasets aiming at evaluating VPR still exists, such as Tokyo 24/7 dataset tokyo247 , ESSEX3IN1 dataset ESSEX3IN1 , INRIA Holidays dataset INRIA , Synthia dataset SYNTHIA , etc.
8.2.2 MRL  for  IPA
To evaluate both VPR and I2P-VPR, it requires the dataset to have both images and LiDAR sweeps. Since I2P-VPR is a newly emerged research, there is very few dataset developed for the evaluation in this field.
KITTI dataset KITTI is a widely used AD benchmark that contains 22 sequences, of which 11 sequences (00-10) present ground-truth trajectories that is originally used for VO or SLAM evaluation. Based on the given pose of each image, ground-truth matches between query and reference image/point cloud can be labelled, and both VPR and I2P-VPR can be evaluated since synchronized LiDAR scans are also provided, we can offline build LiDAR point cloud map using LiDAR SLAM to generate reference point cloud clip as mentioned in Sec. 5.2.
8.2.3 MRL  for  FPE
Some datasets provide ground-truth pose for each image so that we can evaluate the MRL methods aiming at FPE. These dataset usually provide full sensor suite, so we can build visual landmark map (Sec. 4), point cloud map (Sec. 5), HD map (Sec. 6), and even DNN-based implicit map (Sec. 7).
Aachen Day-Night benchmark6dof utilizes reference images to build the scene map, which were taken during daytime with hand-held cameras over a period of about two years. The datasets offers query images taken at daytime and at nighttime so that viewpoint changes and illumination changes are both included. It is popularly used in evaluation for VL-MRL methods.
InLoc Dataset InLoc is designed for large-scale indoor localization that contains significant variation in appearance between queries and the 3D database due to large viewpoint changes, moving furniture, occlusions or changing illumination. The dataset is composed of a database of RGBD images geometrically registered to the floor maps augmented with a separate set of RGB query images taken by hand-held devices to make it suitable for the task of indoor localization. The reference images are 9,972 perspective images consisting of 277 RGBD panoramic images obtained from scanning two buildings at the Washington University in St. Louis. The query images consist of 329 photos using a smart-phone camera annotated with ground-truth 6 DoF camera poses. It is used in evaluation for VL-MRL methods.
KITTI Dataset KITTI provides ground-truth pose for each image, so it can also be used to evaluate MRL for FPE. Since multiple sensor data is provided, it can support evaluation for many kind of MRL.
KAIST Urban KAIST provides LiDAR data and stereo image with various sensors targeting a highly complex urban environment. It captures features in urban environments (e.g. metropolis areas, complex buildings and residential areas). Wen et al. build HD map for it so that it can be used to evaluate HD-MRL methods TM3Loc .
7Scenes posenet is a collection of tracked RGB-D camera frames. The authors used KinectFusion system newcombe2011kinectfusion to obtain the ground truth camera poses, and a dense 3D model. Several sequences were recorded per scene by different users, and split into distinct training and testing sequence sets. It supports the evaluation of almost all the kinds of MRL methods.
Cambridge Landmarks posenet is a large scale outdoor MRL dataset taken around Cambridge University. It contains original video, with extracted image frames labelled with their 6 DOF camera pose and a visual reconstruction of the scene. Similar to 7Scenes posenet , it also supports the evaluation of almost all the kinds of MRL methods.
The mentioned above datasets only cover a small part of dataset for evaluating MRL methods for FPE. Except to these dataset, the famous dataset, CMU Seasons and Extended CMU Seasons benchmark6dof , RobotCar Seasons benchmark6dof , ETH-Microsoft Dataset eth_ms_visloc_2021 is also widely used in evaluation, and Map-free Localization dataset arnold2022mapfree provides a specialised benchmark for evaluating RPR methods when only one reference image is included in the scene map.
8.3 Performance of SOTAs
Here, we provide the performance of some SOTA MRL methods from public publications as a reference to evaluate the advantages and disadvantages of different methods.
8.3.1 MRL  for  IPA
MRL methods for IPA, including VPR and I2P-VPR, tend to retrieve reference images given query image, so they are measured and compared based on the precision and recall metrics of their retrieved results.
VPR: First, we give some performance of VPR methods in public datasets, evaluated with different metrics, e.g., recall@K (in Tab. 2) and maximum recall at 100% precision (in Tab. 3). It shows in Tab. 2 that the differentiable VLAD pooling layer could effectively aggregate dense local descriptors into global features, thus making NetVLAD and its extensions SPE-NetVLAD ; vlaad perform better than average pooling netvlad and GeM pooling GeM in retrieval-based GF-VPR tasks. Then, benefit from the global fusion of features, MixVPR mixvpr achieves improved performance, espeically in challenging Nordland dataset with varying weather and seasons. More recently, pre-trained features from visual foundation model (such as DINOv2 oquab2023dinov2 ) are concluded to be an excellent alternate for GF-VPR, and enable AnyLoc keetha2023anyloc achieves impressive retrieval performance in very different scene. Although GF-VPR methods can fully retrieve reference images, their retrieval accuracy is lower than SOTA LF-VPR methods as shown in Tab. 3. Adding verification with local features yue2 or jointly utilizing local feature and global feature fild2 can achieve pleased accuracy while retrieve most of reference images for each query image.
I2P-VPR: As for cross-modal I2P-VPR between query image and reference point cloud clip, a SOTA performance is achieved by I2P-Rec Li2023I2PRecRI as shown in Fig. 17. With the assistance of mature monocular depth estimation, I2P-Rec Li2023I2PRecRI transform query image and point cloud clip from scene map to BEV images, simplifying the cross-modal retrieval problem and leading to pleased performance. It also concludes such a BEV representation of point cloud can work better than point cloud representation, and better monocular depth estimation leads to better retrieval performance
Table 2: Recall over top-K candidates of different GF-VPR techniques on popular benchmarks.
Method dim Pitts250k-test pittsburgh MSLS-val MSLS SPED SPED Nordland nordland
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
AVG netvlad ${}^{\dagger}$ 2048 62.6 82.7 88.4 59.3 71.9 75.5 54.7 72.5 77.1 4.4 8.4 10.4
GeM GeM ${}^{\dagger}$ 2048 72.3 87.2 91.4 65.1 76.8 81.4 55.0 70.2 76.1 7.4 13.5 16.6
NetVLAD netvlad ${}^{\dagger}$ 32768 86.0 93.2 95.1 59.5 70.4 74.7 71.0 87.1 90.4 4.1 6.6 8.2
AVG netvlad ${}^{\star}$ 2048 78.3 89.8 92.6 73.5 83.9 85.8 58.8 77.3 82.7 15.3 27.4 33.9
GeM GeM ${}^{\star}$ 2048 82.9 92.1 94.3 76.5 85.7 88.2 64.6 79.4 83.5 20.8 33.3 40.0
NetVLAD netvlad ${}^{\star}$ 32768 90.5 96.2 97.4 82.6 89.6 92.0 78.7 88.3 91.4 32.6 47.1 53.3
SPE-NetVLAD SPE-NetVLAD ${}^{\star}$ 163840 89.2 95.3 97.0 78.2 86.8 88.8 73.1 85.5 88.7 25.5 40.1 46.1
Gated NetVLAD vlaad ${}^{\star}$ 32768 89.7 95.9 97.1 82.0 88.9 91.4 75.6 87.1 90.8 34.4 50.4 57.7
CosPlace cosplace ${}^{\star}$ 2048 91.5 96.9 97.9 84.5 90.1 91.8 75.3 85.9 88.6 34.4 49.9 56.5
MixVPR mixvpr ${}^{\star}$ 2048 94.1 98.2 98.8 87.0 92.7 94.2 84.7 92.1 94.4 57.9 73.8 79.0
MixVPR mixvpr ${}^{\star}$ 4096 94.6 98.3 99.0 88.0 92.7 94.6 85.2 92.1 94.6 58.4 74.6 80.0
Method Baidu Mall BaiduMail Gardens Point GardenPoint2 17 Places 17places Pitts-30k pittsburgh St. Lucia RatSLAM Oxford oxford
R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
NetVLAD netvlad 53.10 70.51 58.50 85.00 61.58 77.83 86.08 92.66 57.92 72.95 52.88 74.87
CosPlace cosplace 41.62 55.02 74.00 94.50 61.08 76.11 90.45 95.66 99.59 99.93 91.10 99.48
MixVPR mixvpr 64.44 80.28 91.50 96.00 63.79 78.82 91.52 95.47 99.66 100 90.05 98.95
AnyLoc-GeM-DINOv2 keetha2023anyloc 50.13 70.55 88.00 97.50 63.55 79.56 77.04 87.28 76.91 89.34 81.15 97.38
AnyLoc-VLAD-DINO keetha2023anyloc 61.17 78.32 95.00 98.50 63.79 78.82 83.45 91.99 88.46 94.88 78.53 96.34
AnyLoc-VLAD-DINOv2 keetha2023anyloc 75.22 87.57 95.50 99.50 65.02 80.54 87.66 94.69 96.17 98.84 98.95 100
  • 1
    The results are cited from MixVPR mixvpr and AnyLoc keetha2023anyloc .
  • $\dagger$
    The results reported by the original publications.
  • $\star$
    The variants are trained on GSVcities dataset GSVCities using the same backbone network (ResNet-50 resnet ) with MixVPR mixvpr .
Methodology From Oxford fabmap KITTI KITTI Malaga malaga St. Lucia RatSLAM
CityCentre NewCollege #00 #05 #06 Parking 6L 100909 (12:10) 180809 (15:45)
Image Global Feature-based Visual Place Recognition
NetVLAD netvlad TPAMI’18 0.7148 0.4369 0.9315 0.9108 0.9632 0.3236 0.8046 0.7967
ResNet50-AP-GeM GeM TPAMI’19 0.9056 0.8485 0.9124 0.9470 0.9779 0.6011 0.7926 0.6936
Kazmi et al. Kazmi2019TRO TRO’19 0.7558 0.5109 0.9039 0.8141 0.9739 0.5098 0.8006 0.7255
Image Local Feature-based Visual Place Recognition
FABMAP 2.0 fabmap2 IJRR’11 0.3526 0.6038 0.8222 0.3712 0.6347 n.a. n.a. n.a.
DLoopDetector dbow-tro TRO’12 0.3059 n.a. 0.7243 0.5197 0.8971 0.7475 0.3722 0.3136
Tsintotas et al. AVWs ICRA‘18 0.6595 0.2988 0.9318 0.9420 0.9669 0.8799 0.2627 0.1507
iBoW-LCD ibow-lcd RAL’18 0.8825 0.7940 0.7650 0.5307 0.9553 0.5098 0.7002 0.8750
Tsintotas et al. BoTW RAL’19 n.a. n.a. 0.9750 0.9260 0.8897 0.8500 n.a. n.a.
Yue et al. yue1 IROS’19 0.8984 0.9229 0.9416 0.9494 0.9963 0.8605 n.a. n.a.
Han et al. lcd-pl JFR’21 0.3345 0.5982 0.9145 0.8651 n.a. 0.7251 n.a.
SLCD slcd TII’21 0.4088 0.7529 0.9753 0.8972 0.9313 n.a. n.a. n.a.
Yue et al. yue2 JFR’22 0.9127 0.9463 0.9569 0.9518 0.9963 0.8992 n.a. n.a.
Image Global and Local Feature-based Visual Place Recognition
HTMap htmap TRO’17 0.7968 0.7360 0.9024 0.7588 0.9703 n.a. n.a. n.a.
FILD fild IROS’19 0.6648 0.7674 0.9123 0.8515 0.9338 0.5609 0.7606 0.6696
FILD++ fild2 JFR’22 0.9091 n.a. 0.9492 0.9542 0.9816 0.6274 0.8339 0.8136
Liu et al.liu ICRA’21 0.8601 0.9121 0.9302 0.9253 n.a. n.a. n.a. n.a.
Sequence-based Visual Place Recognition
SeqSLAM seqslam ICRA’12 0.3186 0.5757 0.8758 0.1823 0.6068 n.a. n.a. n.a.
Bampis et al. Bampis2018IJRR IJRR’18 0.4963 0.8087 0.8947 0.8771 0.8015 0.3393 0.6093 0.4979
  • 1
    Most of the results are cited from recent papers yue2 and fild2 .
  • 2
    “n.a.” means the method cannot achieve 100% precision in our implementation with default setups and the result in published papers is unavailable.
[…]
Summary: The objective of MRL methods for IPA is to accurately and completely retrieve reference for each query, which requires robustly measure the similarity between query and reference against shifted viewpoints, changed appearance, dynamic occlusions, and so on. The utilization of pre-trained feature from computer vision community keetha2023anyloc , geometric verification based on local features yue2 , robust representation combined local and global features fild2 , temporal enhancement Bampis2018IJRR may be helpful ways to improve performance for this field. But since VPR and I2P-VPR methods directly use retrieved reference pose as current pose, their localization accuracy is inherently limited, which will be experimental concluded in later paragraphs. So they are generally used as coarse or initial pose approximation in hierarchical localization framework hloc2 .
Additionally, VPR-Bench [^5] https://github.com/MubarizZaffar/VPR-Bench vprbench is an open-sourced VPR evaluation framework with quantifiable viewpoint and illumination invariance, enabling fast and full quantified evaluation for VPR methods. In I2P-VPR field, such public evaluation benchmark is not proposed yet.
Figure 17: Recall over top-K candidates performance of I2P-Rec Li2023I2PRecRI on KITTI dataset KITTI .
8.3.2 MRL  for  FPE
Several MRL methods tend to estimate fine pose. They utilize various kind of scene map and achieve different level of localization accuracy. We first provide performance of RPR, VL-MRL, I2P-MRL, HD-MRL, and NN-MRL respectively. Then, most kind of VL-MRL methods are evaluated on two typical dataset so that their localization accuracy can be fairly compared.
RPR: Towards evaluation of RPR methods, a benchmark called mapfree-reloc-benchmark has been recently proposed  [^6] https://research.nianticlabs.com/mapfree-reloc-benchmark/leaderboard, we suggest readers to refer to the public leader board for detailed information. Up to now, the SOTA solution of RPR is solving pose by decomposing essential matrix between matched features across query image and reference image and then recover absolute scale by a SOTA monocular depth estimation method, DPT DPT . This solution can achieve a localization performance with 1.23m translation error and 11.1° rotation error.
[…]
Day Night
Local Feature Evaluation on Night-time Queries
SIFT sift 5K + NN - 20.4 / 27.6 / 37.8
SIFT + CAPS CAPS + NN - 77.6 / 86.7 / 99.0
SuperPoint superpoint + NN - 73.5 / 79.6 / 88.8
SuperPoint + S2DNet S2DNet - 74.5 / 84.7 / 100.0
SuperPoint + CAPS CAPS + NN - 82.7 / 87.8 / 100.0
D2-Net (SS) d2net 1K + NN - 64.3 / 78.6 / 91.8
D2-Net (SS) d2net 2K + NN - 74.5 / 85.7 / 96.9
D2-Net (SS) d2net 5K + NN - 77.6 / 92.9 / 100.0
D2-Net (MS) d2net 1K + NN - 53.1 / 74.5 / 86.7
D2-Net (MS) d2net 2K + NN - 61.2 / 81.6 / 94.9
R2D2 r2d2 1K + NN - 55.1 / 70.4 / 77.6
R2D2 r2d2 2K + NN - 63.6 / 78.6 / 87.8
R2D2 r2d2 5K + NN - 69.4 / 84.7 / 98.0
ASLFeat aslfeat + NN - 77.6 / 89.8 / 100.0
SEKD (SS) sekd 1K + NN - 30.6 / 33.7 / 38.8
SEKD (SS) sekd 2K + NN - 42.9 / 51.0 / 57.1
SEKD (SS) sekd 5K + NN - 56.1 / 69.4 / 79.6
SEKD (MS) sekd 1K + NN - 35.7 / 42.9 / 50.0
SEKD (MS) sekd 2K + NN - 50.0 / 63.3 / 70.4
SEKD (MS) sekd 5K + NN - 67.3 / 74.5 / 83.7
DISK disk 1K + NN - 60.2 / 72.4 / 81.6
DISK disk 2K + NN - 70.4 / 82.7 / 94.9
DISK disk 5K + NN - 80.6 / 86.7 / 99.0
GLFeat glfeat 5K + NN - 79.6 / 86.7 / 100.0
ALIKE-L alike 1K + NN - 59.2 / 73.5 / 83.7
ALIKE-L alike 2K + NN - 74.5 / 87.8 / 98.0
ALIKED-N(32) aliked 1K + NN - 77.6 / 88.8 / 100.0
ALIKED-N(32) aliked 2K + NN - 76.5 / 87.8 / 100.0
DualRC-Net DRCNet - 79.6 / 88.8 / 100.0
SparseNCNet SparseNCNet - 76.5 / 84.7 / 98.0
Patch2Pix patch2pix - 79.6 / 87.8 / 100.0
Full Localization with HLoc hloc2
SuperPoint superpoint + NN 85.4 / 93.3 / 97.2 75.5 / 86.7 / 92.9
SuperPoint + CAPS CAPS + NN 86.3 / 93.0 / 95.9 83.7 / 90.8 / 96.9
SuperPoint + SuperGlue superglue 89.6 / 95.4 / 98.8 86.7 / 93.9 / 100.0
Patch2Pix patch2pix 84.6 / 92.1 / 96.5 82.7 / 92.9 / 99.0
Patch2Pix (w.CAPS) patch2pix 86.7 / 93.7 / 96.7 85.7 / 92.9 / 99.0
Patch2Pix (w.SuperGlue) patch2pix 89.2 / 95.5 / 98.5 87.8 / 94.9 / 100.0
  • 1
    Some results are cited from Patch2Pix patch2pix and we do not know the limitation of number of extracted features or matched features. The results with limited number of features are denotes like “SIFT 5K” and cited from ALIKED aliked and GLFeat glfeat .
Table 4: Localization success ratio of some VL-MRL methods on Aachen Day-Night Benchmark (v1.0) benchmark6dof .
Table 5: Localization success ratio of some VL-MRL methods on InLoc Benchmark  InLoc .
[…]
DUC1 DUC2
SuperPoint superpoint + NN 40.4 / 58.1 / 69.7 42.0 / 58.8 / 69.5
D2Net d2net + NN 38.4 / 56.1 / 71.2 37.4 / 55.0 / 64.9
R2D2 r2d2 + NN 36.4 / 57.6 / 74.2 45.0 / 60.3 / 67.9
SuperPoint + SuperGlue superglue 49.0 / 68.7 / 80.8 53.4 / 77.1 / 82.4
SuperPoint + CAPS CAPS + NN 40.9 / 60.6 / 72.7 43.5 / 58.8 / 68.7
SIFT + CAPS CAPS + NN 38.4 / 56.6 / 70.7 35.1 / 48.9 / 58.8
SparseNCNet SparseNCNet 41.9 / 62.1 / 72.7 35.1 / 48.1 / 55.0
Patch2Pix patch2pix 44.4 / 66.7 / 78.3 49.6 / 64.9 / 72.5
Patch2Pix (w.SuperPoint+CAPS) patch2pix 42.4 / 62.6 / 76.3 43.5 / 61.1 / 71.0
Patch2Pix (w.SuperGlue) patch2pix 50.0 / 68.2 / 81.8 57.3 / 77.9 / 80.2
Table 6: 6 DoF localization errors of some I2P-MRL methods on KITTI Odometry Dataset KITTI
Method Mean Median
Rot. (${}^{\circ}$$\downarrow$ Transl. (m) $\downarrow$ Rot. (${}^{\circ}$$\downarrow$ Transl. (m) $\downarrow$
Geometry-based Cross-modal Localization
Caselitz et al. mono_i2p_vo 1.65 0.30 n.a. n.a.
Cross-modal Relative Pose Regression
CMRNet cmrnet 1.09 0.26 1.39 0.51
HyperMap hypermap n.a. n.a. 1.42 0.48
POET miao2023poses 0.91 0.25 0.79 0.20
I2PNet I2PNet 0.74 0.08 0.67 0.07
Cross-modal Matching-based Localization
CMRNet++ cmrnet++ n.a. n.a. 1.46 0.55
I2D-Loc I2DLoc n.a. n.a. 0.70 0.18
  • 1
    The measured 6 DoF localization errors include 3 DoF rotation errors (Rot.) and 3 DoF translation errors (Transl.).
VL-MRL: As for VL-MRL methods, as shown in Tab. 4 and Tab. 5, many recent SOTA method in local feature extraction superpoint ; disk ; CAPS ; aslfeat and matching superglue ; DRCNet ; SparseNCNet can be readily integrated into HLoc framework as alternates to originally used HFNet hloc2 and NN matching. We report a widely used evaluation metric here, which is the ratio of successfully localized queries within three defined tolerances of localization errors. If there were enough correctly matched local features between query image and visual landmark map, HLoc can achieve an excellent localization performance. From Tab. 4, it can be seen that extracting more local features contribute to localization, but MS detection is not constantly positive for localization d2net . The same key point detectors with more advanced descriptors will generate more accurate matches so they can achieve better performance S2DNet ; CAPS .
From Tab. 4 and 5, we can also see that accurate localization in indoor scene InLoc is much more difficult than localization in urban scene benchmark6dof due to lack of textures. In current framework of VL-MRL methods that local features are widely used as scene representation, textures of scene still be necessary.
I2P-MRL: Tab. 6 shows some cross-modal MRL methods on KITTI 00 sequence KITTI . The learning-based I2P-MRL methods cmrnet ; cmrnet++ ; hypermap ; miao2023poses ; I2PNet ; I2DLoc seem to outperform geometry-based one mono_i2p_vo with a lot of margin. The SOTA I2P-MRL methods can achieve an impressive localization performance with 0.08 m in translation error and 0.74°in rotation error I2PNet . However, these I2P-MRL methods are trained and tested on human-defined data that the displacement between query image and reference point cloud is randomly determined within a defined threshold and it is often limited. So, their performance in practical applications is still waiting for evaluation.
Table 7: 3 DoF Localization errors of some HD-MRL methods.
Method Dataset Lon. (m) $\downarrow$ Lat. (m) $\downarrow$ Heading (°) $\downarrow$
Traditional methods
Choi et al. Choi2019Access 16km highways 0.12 0.18 n.a.
21km highways 0.10 0.25 n.a.
Xiao et al. Xiao2018TISC KITTI 04 KITTI 0.345
Liao et al. Liao2019CoarseToFineVL KAIST 26 KAIST 0.64 0.90
Virtual KITTI gaidon2016virtual 0.29 0.32
MLVHM MLVHM Campus 0.21
Urban 0.29
Wen et al. Wen2020IV Campus 0.14
Urban 0.25
ROADMap RoadMap 22km Urban 0.04 0.04 0.12
Wang et al. Wang2021ICRA Urban canyons 0.46 0.13 0.11
Viaducts 0.45 0.13 0.11
[…]
KAIST 23 KAIST 0.32 0.05 0.61
KAIST 26 KAIST 0.16 0.05 0.48
3.2km Shougang Park 0.16 0.07 2.87
2.0km Shougang Park 0.23 0.12 2.76
1.3km Shougang Park 0.18 0.06 2.05
Learning-based methods
Zhang et al. Zhang2023TIV Highway 0.17 0.04 n.a.
BEV-Locator bevlocator nuScenes nuscenes2019 0.18 0.08 0.51
Qcraft 0.14 0.05 0.25
EgoVM he2023egovm nuScenes nuscenes2019 0.11 0.03 0.09
self-collected dataset 0.04 0.03 0.08
  • 1
    The measured 3 DoF localization errors include longitudinal (Lon.), lateral (Lat.), and Heading angle errors.
HD-MRL: In Tab. 7, we collect the results of some HD-MRL methods from their publications. Since the public HD map data is very few in existing dataset, researchers often collect and build their own dataset to evaluate the proposed methods. For fair and full comparison, we also list the evaluated scene/dataset for each method. It can be seen that the tightly coupling of VO and HD-MRL TM3Loc improves the accuracy in long-term localization. The learning-based methods effectively solve the data association problem between 2D images and HD map elements, thus outperforming traditional methods, and they can localize without the requirements of online VO threads. BEV-Locator he2023egovm estimates pose by regression whereas EgoVM he2023egovm relies on exhaustively searching pose candidates, the latter solution achieves better performance. It should be noticed that HD map is a very compact format of map representation. In MLVHM MLVHM , the HD maps stored in ASCII format only requires about 50 KB storage space per kilometer, compared to about 600 MB per kilometer of the raw point cloud map. It is not the only case that the map size of EgoVM is 0.35MB/km compared to 5.92MB/km of DA4AD DA4AD . Therefore, the compactness and accuracy of HD map makes it a popular choice for localization in AD.
Table 8: Localization results on Cambridge Landmarks dataset posenet .
Method College Hospital Shops St. Mary’s
Localization with Geo-tagged Frame Map
Visual Place Recognition
VLAD tokyo247 2.80,5.7° 4.01,7.1° 1.11,7.6° 2.31,8.0°
VLAD+Inter Sattler2019UnderstandingTL 1.48,4.5° 2.68,4.6° 0.90,4.3° 1.62,6.1°
Relative Pose Regression
EssNet EssNet 0.76,1.9° 1.39,2.8° 0.84,4.3° 1.32,4.7°
NC-EssNet EssNet 0.61,1.6° 0.95,2.7° 0.70,3.4° 1.12,3.6°
RelocGNNDSAC2 0.48,1.0° 1.14,2.5° 0.48,2.5° 1.52,3.2°
Localization with Visual Landmark Map
ActiveSearch VPS 0.42,0.6° 0.44,1.0° 0.12,0.40° 0.19,0.5°
InLocInLoc 0.18,0.6° 1.2,0.6° 0.48,1.0° 0.46,0.8°
Localization with Point Cloud Map
PixLoc pixloc 0.14/0.2° 0.16/0.3° 0.05/0.2° 0.10/0.3°
Localization with Learnt Implicit Map
Single Scene Absolute Pose Regression
PoseNet posenet 1.92,5.40° 2.31,5.38° 1.46,8.08° 2.65,8.48°
BayesianPN BayesianPN 1.74,4.06° 2.57,5.14 ° 1.25,7.54° 2.11,8.38°
LSTM-PN LSTMPN 0.99,3.65° 1.51,4.29° 1.18,7.44° 1.52,6.68°
SVS-Pose SVS-Pose 1.06,2.81° 1.50,4.03° 0.63,5.73° 2.11,8.11°
GPoseNet Cai2018AHP 1.61,2.29° 2.62,3.89° 1.14,5.73° 2.93,6.46°
PoseNetLearn PoseNet2 0.99,1.06° 2.17,2.94 ° 1.05,3.97° 1.49,3.43°
GeoPoseNet PoseNet2 0.88,1.04° 3.20,3.29° 0.88,3.78° 1.57,3.32°
MapNet mapnet 1.07,1.89° 1.94,3.91° 1.49,4.22° 2.00,4.53°
IRPNet Shavit2020DoWR 1.18,2.19° 1.87,3.38° 0.72,3.47° 1.87,4.94°
Multiple Scene Absolute Pose Regression
MSPN MSPN 1.73,3.65° 2.55,4.05° 2.92,7.49° 2.67/6.18°
MS-TransMS-Trans1 0.83,1.47° 1.81,2.39° 0.86,3.07° 1.62,3.99°
c2f-MS-Trans MS-Transformer2 0.70,2.69° 1.48,2.94 ° 0.59,2.88° 1.14,3.88°
Scene Coordinate Regression
DSAC DSAC 0.30,0.5° 0.33,0.6° 0.09,0.4° 0.55,1.6°
DSAC++ Less_is_More 0.18,0.3° 0.20,0.3° 0.06,0.3° 0.13,0.4°
$*{}^{*}$ DSAC2 0.18,0.3° 0.21,0.4° 0.05,0.3° 0.15,0.5°
SANet SANet 0.32/0.5° 0.32/0.5° 0.10/0.5° 0.16/0.6°
Absolute Pose Regression with NeRF Enhancement
DFNet DFNet n.a. 0.46,0.87° 0.16,0.59° 0.50,1.49°
LENS Moreau2021LENSLE 0.33,0.5° 0.44,0.9° 0.27,1.6° 0.53,1.6°
NeRF as Pose Estimator
NeRF-Loc liu2023nerfloc 0.11,0.2° 0.18,0.4° 0.04,0.2° 0.07,0.2°
Table 9: Localization results on 7Scenes dataset posenet .
Chess Fire Heads Office Pumpkin Kitchen Stairs Average
Localization with Geo-tagged Frame Map
Visual Place Recognition
VLAD tokyo247 0.21,12.5° 0.33,13.8° 0.15,14.9° 0.28,11.2° 0.31,11.2° 0.30,11.3° 0.25,12.3° 0.26,12.46°
VLAD+InterSattler2019UnderstandingTL 0.18,10.0° 0.33,12.4° 0.14,14.3° 0.25,10.1° 0.26,9.4° 0.27,11.1° 0.24,14.7° 0.24,11.71°
Relative Pose Regression
NN-Net Zakaria2017ICCVW 0.13,6.5° 0.26,12.7° 0.14,12.3° 0.21,7.4° 0.24,6.4° 0.24,8.0° 0.27,11.8° 0.21,9.30°
RelocNetRelocNet 0.12,4.1° 0.26,10.4° 0.14,10.5° 0.18,5.3° 0.26,4.2° 0.23,5.1° 0.28,7.5° 0.21,6.73°
EssNet EssNet 0.13,5.1° 0.27,10.1° 0.15,9.9° 0.21,6.9° 0.22,6.1° 0.23,6.9° 0.32,11.2° 0.22,8.03°
NC-EssNet EssNet 0.12,5.6° 0.26,9.6° 0.14,10.7° 0.20,6.7° 0.22,5.7° 0.22,6.3° 0.31,7.9° 0.21,7.50°
RelocGNNDSAC2 0.08,2.7° 0.21,7.5° 0.13,8.70° 0.15,4.1° 0.15,3.5° 0.19,3.7° 0.22,6.5° 0.16,5.24°
CamNet CamNet 0.04,1.73° 0.03, 1.74° 0.05,1.98° 0.04,1.62° 0.04,1.64° 0.04,1.63° 0.04,1.51° 0.04,1.69°
Localization with Visual Landmark Map
Active Search VPS 0.04,2.0° 0.03,1.5° 0.02,1.5° 0.09,3.6° 0.08,3.1° 0.07,3.4 ° 0.03,2.2° 0.05,2.47°
InLocInLoc 0.03,1.05° 0.03,1.07° 0.02,1.16° 0.03,1.05° 0.05,1.55° 0.04,1.31° 0.09,2.47° 0.04,1.44°
Localization with Point Cloud Map
PixLoc pixloc 0.02/0.8° 0.02/0.7° 0.01/0.8° 0.03/0.8° 0.04/1.2° 0.03/1.2° 0.05/1.3° n.a.
Localization with Learnt Implicit Map
Single Scene Absolute Pose Regression
PoseNet posenet 0.32,8.12° 0.47,14.4° 0.29,12.0° 0.48,7.68° 0.47,8.42° 0.59,8.64° 0.47,13.8° 0.44,10.44°
BayesianPN BayesianPN 0.37,7.24° 0.43,13.7° 0.31,12.0° 0.48,8.04° 0.61,7.08° 0.58,7.54° 0.48,13.1° 0.47,9.81°
LSTM-PN LSTMPN 0.24,5.77° 0.34,11.9° 0.21,13.7° 0.30,8.08° 0.33,7.00° 0.37,8.83° 0.40,13.7° 0.31,9.85°
GPoseNet Cai2018AHP 0.20,7.11° 0.38,12.3° 0.21,13.8° 0.28,8.83° 0.37,6.94° 0.35,8.15° 0.37,12.5° 0.31,9.95°
PoseNetLearnPoseNet2 0.14,4.50° 0.27,11.8° 0.18,12.1° 0.20,5.77° 0.25,4.82° 0.24,5.52° 0.37,10.6° 0.24,7.87°
GeoPoseNetPoseNet2 0.13,4.48° 0.27,11.3° 0.17,13.0° 0.19,5.55° 0.26,4.75° 0.23,5.35 ° 0.35,12.4° 0.23,8.12°
IRPNetShavit2020DoWR 0.13,5.64° 0.25,9.67° 0.15,13.1° 0.24,6.33° 0.22,5.78° 0.30,7.29° 0.34,11.6° 0.23,8.49°
AtLocWang2019AtLocAG 0.10,4.07° 0.25,11.4° 0.16,11.8° 0.17,5.34° 0.21,4.37 ° 0.23,5.42° 0.26,10.5° 0.20,7.56°
Multiple Scene Absolute Pose Regression
MSPNMSPN 0.09/4.76° 0.29/10.5° 0.16/13.1° 0.16/6.80° 0.19/5.50° 0.21/6.61° 0.31/11.6° 0.20/7.56°
MS-TransMS-Trans1 0.11,4.66° 0.24,9.60 ° 0.14,12.2° 0.17,5.66° 0.18,4.44° 0.17,5.94 ° 0.26,8.45° 0.18,7.28°
c2f-MS-Trans MS-Transformer2 0.10,4.60° 0.24,9.88° 0.12,12.3° 0.16,5.64° 0.16,4.42° 0.16,6.39° 0.25,7.76° 0.17/7.28°
Scene Coordinate Regression
DSAC DSAC 0.02,0.7° 0.03,1.0° 0.02,1.30° 0.03,1.0° 0.05,1.30° 0.05,1.5° 1.90,49.4° 0.30,8.03°
DSAC++Less_is_More 0.02,0.5° 0.02,0.9° 0.01,0.8 0.03,0.7° 0.04,1.1° 0.04,1.1° 0.09,2.6° 0.04,1.10°
$*{}^{*}$ DSAC2 0.02,1.1° 0.02,1.0° 0.01,1.8° 0.03,1.2° 0.04,1.4° 0.03,1.7° 0.04,1.4° 0.03,1.37°
SANet SANet 0.03/0.9° 0.03/1.1° 0.02/1.5° 0.03/1.0° 0.05/1.3° 0.04/1.4° 0.16/4.6° n.a.
Absolute Pose Regression with NeRF Enhancement
DFNet DFNet 0.04,1.48° 0.04,2.16° 0.03,1.82° 0.07,2.01° 0.09,2.26° 0.09,2.42° 0.14,3.31° 0.07,2.21°
Direct-PN+U DPN 0.09,2.77° 0.16,4.87° 0.10,6.64° 0.17,5.04° 0.19,3.59° 0.19,4.79° 0.24,8.52° 0.16,5.17°
LENS Moreau2021LENSLE 0.03,1.3° 0.10,3.7° 0.07,5.8° 0.07,1.9° 0.08,2.2° 0.09,2.2° 0.14,3.6° 0.08,3.0°
NeRF as Pose Estimator
NeRF-Loc liu2023nerfloc 0.02/1.1° 0.02/1.1° 0.01/1.9° 0.02/1.1° 0.03/1.3° 0.03/1.5° 0.03/1.3° n.a.
NN-MRL: Then, we discuss about MRL with learnt implicit map in Sec. 7. According to the localization results on medium-scale Cambridge Landmarks dataset posenet in Tab. 8 and small-scale 7Scenes dataset posenet in Tab. 9, we find that the localization accuracy of APR methods is not pleased, and sometimes they even perform worse than VPR methods that only give an approximated pose. APR performs much worse than VL-MRL methods. When enhanced by NeRF, the performances of APR methods are significantly improved, and achieve comparable performances with VL-MRL methods. As a surprise, we find that SCR methods work well and achieve impressive localization accuracy in both medium-scale outdoor and small-scale indoor scenes. DSAC DSAC and its extended works Less_is_More ; DSAC2 even work better than VL-MRL methods in some cases. SCR may be a promising trend in MRL research in the future. But its heavy cost of training will block its wide usage. As for NeRF, it can significantly boost APR methods DFNet ; DPN ; Moreau2021LENSLE and it also can work well as a pose estimator liu2023nerfloc . We are still waiting for the usage of NeRF in large-scale and long-term localization tasks since building NeRF for unbounded scene is a much more challenging problem compared to indoor scene.
Summary: Finally, we compare many MRL methods with various kind of scene map on outdoor Cambridge Landmarks posenet and indoor 7Scenes posenet , as listed in Tab. 8 and Tab. 9. In outdoor scene, the rankings of MRL methods with regard to localization accuracy should be: NeRF pose estimator liu2023nerfloc $>$ I2P-MRL pixloc $>$ SCR $>$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. And in indoor scene, the rankings of MRL methods are slightly changed as: NeRF pose estimator liu2023nerfloc $\approx$ I2P-MRL pixloc $\approx$ SCR $\approx$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. It can be seen that in small-scale indoor scene, although APR and RPR work worse than other MRL methods, they can also achieve pleased localization performance. Given the low requirement of built map in APR and RPR, they can be used in some applications that do not need high-precision localization, such as VR. VPR only retrieve visually similar reference image and regard the reference pose as an approximated query pose, so its localization accuracy is very limited when reference images are few in outdoor scene. In indoor scene, most MRL methods with various kinds of scene map achieve comparable localization performance. But in outdoor scene, the gaps among different methods are more significant. Using NeRF as pose estimator like NeRF-Loc liu2023nerfloc , I2P-MRL pixloc , SCR, and APR+NeRF seem to be promising ways in MRL researches. Adapting NeRF-based localization to large-scale unbounded scenes, reducing the requirement of high-precision point cloud map for I2P-MRL, reducing the training cost of SCR should be explored to make these MRL methods more practical.
9 Opening Discussion
In this section, we discuss about some opening questions in MRL field, and provide our opinion. The answers of following questions are still unavailable, we only hope our personal opinion can motivate further contributions in MRL research.
9.1 Are Explicit 3D Models Necessary in Visual Localization?
Traditional MRL methods require an explicit 3D scene map, e.g., visual landmarks map in VL-MRL, point cloud map in I2P-MRL, and vectorized map in HD-MRL. An explicit 3D model of the scene appears to be necessary in the early stage of MRL research. However, with the recent advancements in MRL, some techniques no longer require an explicit 3D model, such as VPR, RPR, and NN-MRL. Therefore, a natural question arises: are explicit 3D models necessary for accurate visual localization? This is an extension of a widely known question: are large-scale 3D models really necessary for accurate visual localization?3dmodel .
As shown in Tab. 8 and 9, it can be seen that some 2D image-based methods without explicit map, namely, VPR and RPR, perform worse than 3D model-based method with explicit map such as VL-MRL and I2P-MRL. However, 2D image-based methods with a learnt implicit map, i.e., NN-MRL, can achieve comparable or even better performances. Specially, NeRF-Loc liu2023nerfloc has been found to localize with centimeter-level accuracy, surpassing VL-MRL and I2P-MRL by a considerable margin. Thus, it is believed that the learnable implicit map has the potential to be a superior alternative to the explicit map for accurate localisation. The explicit 3D map is not strictly necessary in current deep learning era.
9.2 How to Select Specific Localization Method?
Hundreds of MRL researches are proposed each year, relying on various kinds of scene map and achieving differing levels of localization accuracy. Therefore, one may consider how to choose the appropriate MRL methods for a specific autonomous task?
As shown in the experimental results in Sec. 8.3, the utilized scene map of MRL usually determines the localization accuracy that a particular kind of MRL can achieve. 1) VPR methods estimate current pose by retrieving historical posed images. Their localization accuracy is limited, but they solely require a 2D image map and can obtain reference images with overlapped view-of-field. As a result, they are often utilized as a coarse step in hierarchical localization framework hloc1 ; hloc2 to acquire co-visible query-reference image pairs. 2) With a slightly better localization accuracy, RPR and APR methods can be employed in some applications that do not necessitate high-precision poses but call for a low-cost and light scene map. 3) VL-MRL and I2P-MRL methods require an offline mapping stage to obtain 3D scene map, i.e., visual landmarks map or point cloud map, so their cost of mapping is significantly higher than previously mentioned methods. With these precise and explicit 3D scene map, these two kind of methods achieve outstanding localization performance under most circumstances. However, when aiming for autonomy in large-scale scenarios, these two kinds of scene map require significant storage space, making them difficult to deploy on resource-limited platform, like intelligent vehicles. 4) HD map is the most light and compact form of scene map, so HD-MRL methods are widely applied in autonomous driving task. But HD Map is limited to human-made scenes with semantic map elements, and it is not suitable for natural scenes. 5) Recently developed NN-MRL methods utilize DNN model to represent scenes, which have demonstrated impressive success in small-scale indoor scene and medium-scale urban scenes. However, the efficacy of NN-MRL has yet to be evaluated in unbounded scenes. For instance, SCR methods require dense ground-truth data of scene coordinates, which is hard to be obtained in large-scale scenes. Training NeRF for unbounded dynamic scenes remains a formidable task, rendering MRL based on NeRF currently unsolvable.
In summary, we should choose specific MRL methods based on its required localization accuracy, supported platform, and types of applied scenarios (including textures, scene size, etc.).
9.3 What’s Next for Visual Localization?
Developed over serveral decades, MRL continues to attract growing attention from both academia and industry. We would like to discuss about an interesting and instructive problem: ” What are the future trends of visual localization research?”
End-to-end pipeline: The robustness against environmental changes should be well considered by MRL methods for long-term autonomy. The current query image often has different visual conditions with reference data stored in the scene map. Most of existing solutions rely on image feature algorithms sift ; superpoint to attain condition-invariant image representation. However, these features often fail to achieve comparable performance to that on feature matching tasks. For instance, the matching accuracy of D2-Net d2net is inferior to SuperPoint superpoint on HPatches benchmark hpatches1 ; hpatches2 , but whtn it comes to localization benchmark benchmark6dof , D2-Net d2net yields better results than SuperPoint superpoint . This discrepancy between feature matching and localization might be attributed to the fact that human-defined supervision of features. Some recent approaches suggest building the whole localization pipeline into an E2E manner, incorporating feature extraction, matching, and pose estimation, and training such localization models by supervising them on estimated poses I2PNet ; I2DLoc ; Stumberg2019GNNetTG ; pixloc . Such pose supervision can facilitate the learning of involved geometry priors in MRL and enhance the global consistency of feature matching. This kind of image feature are specifically learned for MRL, enabling the closure of the gap between feature matching and localization, thus becoming a more reasonable solution for representation in MRL research. It is pertinent to note that HD-MRL is encountered with similar challenges alongside feature-based localization. Since semantic instance-level matching is sparse and difficult, some E2E solutions prefer to make models extract and match high-dimensional features for both images and vectorized map elements, and then estimate poses, leading to improved localization performances bevlocator ; he2023egovm . We believe E2E localization system will be a promising trend.
Resource-friendly MRL: To achieve real-time ego-pose estimation in large-scale scenarios during long-term exploration, we must take into account the storage demands of scene map and the running efficiency of localization algorithms. For example, high-level autonomous vehicles in unbounded driving areas generally depend on vectorized HD map, rather than visual landmarks map and point cloud map that are commonly used in robotic applications within confined scenes roadsidemapping . In the field of robotic localization, several recommendations are available to lighten scene map, e.g., SceneSqueezer SceneSqueezer , NeuMap neumap , and HyperMap hypermap , which significantly decrease the map size while ensuring localization accuracy. There is no doubt that MRL will prefer a lightweight and compact scene map in the future. Aiming at high efficiency, MRL methods are commonly built in coarse-to-fine framework where VPR is applied to coarsely estimate poses hloc1 ; hloc2 . Recent advances in MRL methods enable us to employ NN-MRL solutions as a coarse step, such as APR in LATITUDE LATITUDE . Additionally, improving the efficiency of key components in practical MRL pipelines is also essential. ALIKE alike efficiently extracts local features while LightGLue lindenberger2023lightglue achieves high-speed matching. Efforts to reduce computational cost of MRL methods should be encouraged, as this will stimulate the proposal of a growing number of practical MRL solutions for resource-limited platforms.
MRL with new kind of map: In recent years, numerous MRL methods that use new types of scene map have been created. For example, MeshLoc meshloc employs a dense mesh-based map rather than visual landmarks map and achieves superior performance. Similarly, NN-MRL methods utilize implicit neural network-based scene map as a substitute for traditional explicit 3D structure-based scene map, e.g., visual landmarks map and point cloud map. Proposing new representation formats of scene map is still an area of research that requires further exploration in MRL. The ideal scene map should be lightweight, compact, and easy to build and deploy. Moreover, it ought to provide comprehensive information for autonomy. The reference data for localization in the scene map should be sufficiently robust against potential changes in actual scene. Current scene map is challenging to accomplish all of these criteria, so we believe it is necessary to develop new kind of scene map for MRL, which will bring about impressive improvement to MRL researches.
[…]
10 Conclusions
In this survey, we formulate MRL method as a interaction procedure between query image and scene map where poses are estimated, and then we systemically review MRL methods based on the representation format of utilized scene map, that is MRL using geo-tagged frame map, visual landmark map, point cloud map, HD map, and more recently proposed learnt implicit map. Each kind of MRL methods and the related component of them are fully reviewed. Also, we provide a review about evaluation of MRL methods, and draw some conclusions based on the evaluation results of typical algorithms. We provide some opening problem in this field and give our personal opinions. Finally, as a continuous contribution to the community, we list the reviewed papers and datasets on the website so that researches can readily find the best-matched MRL method based on their interests.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants U22A20104 and 52372414, and Beijing Municipal Science and Technology Commission (Grant No.Z221100008122011).